### v1

# Collaborative Perception Datasets for Autonomous Driving: A Review

Abstract
Collaborative perception has attracted growing interest from academia and industry due to its potential to enhance perception accuracy, safety, and robustness in autonomous driving through multi-agent information fusion. With the advancement of Vehicle-to-Everything (V2X) communication, numerous collaborative perception datasets have emerged, varying in cooperation paradigms, sensor configurations, data sources, and application scenarios. However, the absence of systematic summarization and comparative analysis hinders effective resource utilization and standardization of model evaluation. As the first comprehensive review focused on collaborative perception datasets, this work reviews and compares existing resources from a multi-dimensional perspective. We categorize datasets based on cooperation paradigms, examine their data sources and scenarios, and analyze sensor modalities and supported tasks. A detailed comparative analysis is conducted across multiple dimensions. We also outline key challenges and future directions, including dataset scalability, diversity, domain adaptation, standardization, privacy, and the integration of large language models. To support ongoing research, we provide a continuously updated online repository of collaborative perception datasets and related literature: https://github.com/frankwnb/Collaborative-Perception-Datasets-for-Autonomous-Driving.
Index Terms:
Autonomous Driving, Collaborative Perception, Datasets, V2X Communication, Perception Tasks
I Introduction
Perception serves as a foundational component of Autonomous Driving (AD) systems [1], [2], [3], [4], [5], [6], [7], [8], [9], enabling the vehicle to perceive its surroundings through various onboard sensors and to perform real-time tasks such as object detection [10][11], semantic segmentation, tracking, and trajectory prediction in dynamic traffic environments [12]. The precision and reliability of the perception system directly impact the safety and decision-making capability of autonomous vehicles, making it a critical aspect to achieve high-level AD. In recent years, significant progress has been made in single-vehicle perception, driven by progress in computer vision [13][14][15], deep learning [16][17], and sensor technology [18][19]. Numerous studies have leveraged cameras [20][21][22], LiDAR [23][24], Radar [25][26], and multi-sensor fusion [27][28][29][30][31][32] to enhance environmental perception in AD systems. However, traditional single-vehicle perception is constrained by limited fields of view, sensor blind spots, and environmental occlusions, making it insufficient to meet the high-reliability requirements of AD in complex and dynamic traffic environments.
Figure 1: Development Trends of Collaborative Perception Datasets.
To overcome the limitations of single-vehicle perception, Collaborative Perception (CP) leverages Vehicle-to-Everything (V2X) communication to enable real-time information sharing among vehicles and infrastructure [33], thereby enhancing the perception of occluded and long-range regions, improving driving safety, and expanding the Operational Design Domain (ODD) [34][35]. For instance, at an urban intersection, a single vehicle may fail to detect pedestrians or oncoming vehicles that are occluded by buildings or other obstacles. By leveraging V2X connectivity, autonomous vehicles can receive supplementary perception data from roadside infrastructure or neighboring vehicles, thereby improving situational awareness and reducing accident risk. V2X communication encompasses Vehicle-to-Vehicle (V2V), Vehicle-to-Infrastructure (V2I), Infrastructure-to-Infrastructure (I2I), and Vehicle-to-Network (V2N) links, empowering AD systems to overcome the inherent limitations of individual perception and achieve a more holistic understanding of their surroundings.
High-quality datasets are essential for advancing algorithm development and performance evaluation in CP research [36]. Existing single-vehicle perception datasets, such as KITTI [37], nuScenes [38], and Waymo [2], are primarily designed for individual perception tasks and do not support multi-vehicle or infrastructure-level information exchange, rendering them unsuitable for direct use in CP studies.
In recent years, numerous CP datasets, as illustrated in Fig. 1, have been released to support various scenarios, tasks, and sensor modalities, including DAIR-V2X [39], OPV2V [40], V2X-Sim [41], and V2X-Seq [42]. While these datasets offer valuable resources for the CP community, they differ significantly in terms of scale, scene diversity, annotation quality, sensor configurations, and the simulation of communication latency. A systematic review and analysis of existing CP datasets can help researchers better understand their characteristics, application boundaries, and limitations, thereby providing scientific guidance for future investigations. Furthermore, identifying existing gaps and prospective research directions will contribute to the continued advancement of CP technologies.
I-A Roadmap
Figure 2: Chronological overview of collaborative perception datasets, including roadside perception, V2V, V2I, V2X, and I2I datasets.
Fig. 2 illustrates the chronological evolution of CP datasets. Overall, the development of these datasets has progressed through several key phases, each marking a significant shift in perception paradigms or improvements in dataset diversity. Changes in collaboration models, increasing scene complexity, the expansion of supported tasks, and the continuous growth in sensor configurations have all contributed to the advancement of CP technologies.
In the early stages, datasets were primarily focused on infrastructure-based perception, as exemplified by Ko-PER [43], CityFlow [44], and the INTERACTION [45]. However, these datasets did not support V2V or V2I information exchange, thereby limiting their applicability to CP research. With the advancement of V2X communication technologies, the number of related datasets has steadily increased. Around 2020, several V2V and V2I datasets were released, including T & J [46], V2V-Sim [47], CoopInf [48], and CODD [49], which collectively accelerated progress in CP research. However, datasets from this period still faced limitations in terms of scene complexity, sensor modality diversity, and task coverage. As the field progressed, newer CP datasets began to target a broader range of scenarios. For example, Adver-City [50] focuses on adverse weather conditions, DeepAccident [51] addresses accident prediction, and H-V2X [52] is tailored for highway environments. These datasets have significantly enhanced the adaptability of CP systems.
Since 2022, CP dataset development has increasingly emphasized multi-modal fusion and real-world deployment. Datasets such as V2V4Real [53], V2X-Real [54], and DAIR-V2X-C [39] incorporate real-world data collection to overcome the limitations of synthetic datasets in terms of authenticity and distribution. Simultaneously, newer datasets demonstrate broader task coverage and richer sensor modalities. For instance, V2X-Seq [42], V2XReID [55], and DAIR-V2X-Traj [56] support a variety of perception tasks, while V2X-Radar [57] and V2X-R [58] incorporate 4D radar to enhance perception quality. Moreover, the emergence of I2I datasets such as InScope [59] and Rcooper [60] has further enriched the collaborative paradigms within CP. In general, the evolution of CP datasets reflects a continuous trend toward greater complexity, richer modalities, and more specialized scenarios.
I-B Comparison to Related Reviews
We aim to clarify the distinctions between this review and previous related reviews [61], [36], [62], [63], [64], [65], [66], [67], [68], [69], [70]. While some existing reviews partially overlap with ours in terms of dataset coverage [71] [72], they omit several important CP datasets and recent benchmark methodologies introduced in the past few years. Furthermore, their classification frameworks tend to be narrow and often lack systematic summarization and in-depth comparative analysis. In addition, some reviews concentrate primarily on vehicle-centric datasets [61][36][62][63], or explore specific aspects of CP such as multimodal fusion [64][65] or certain CP techniques [66][67][68][69][70], without offering a comprehensive overview of dataset evolution and their applicability across various scenarios. In contrast, this review presents the most up-to-date and complete collection of CP datasets, encompassing a broader spectrum of perspectives. It also delivers an in-depth discussion on dataset diversity, key challenges, and emerging research directions.
Figure 3: Organization of this paper.
I-C Main Contributions
In this review, we present a systematic and comprehensive analysis of CP datasets for AD, as summarized in Fig. 3. The main contributions of this work are as follows:
Comprehensive Review of CP Datasets: To the best of our knowledge, this is the first dedicated and comprehensive review focused specifically on CP datasets. We systematically categorize and analyze available datasets based on collaboration types, data sources, sensor modalities, and application scenarios.
Task-Oriented Dataset Classification: We classify CP datasets according to core perception tasks in AD, including 3D object detection, semantic segmentation, and object tracking and prediction. This task-centric taxonomy facilitates effective dataset selection based on research needs. In addition, we cover the latest developments in CP datasets, providing a timely and up-to-date overview.
Dataset Comparison and Emerging Trends: We conduct multi-dimensional comparisons, highlight high-impact datasets, and discuss key challenges and future directions such as standardization, LLM-driven annotation, and data diversity.
I-D Review Structure
The structure of this paper is illustrated in Fig. 3. To better present our review, the remainder of this paper is organized as follows. Section II introduces datasets under different collaborative paradigms. Section III compares the characteristics of simulated and real-world datasets. Section IV reviews existing CP datasets from the perspective of different application scenarios. Section V analyzes datasets based on sensor modalities. Section VI summarizes and analyzes the CP tasks supported by these datasets, along with their corresponding evaluation metrics. Section VII provides a comprehensive analysis of existing CP datasets, focusing on their availability, quality, and impact. Section VIII discusses the challenges and trends in CP datasets. Section IX concludes the paper and outlines future research directions. A summary of all abbreviations used throughout the paper is provided in Table I.
Table I: Summary of commonly used abbreviations in this review.
Abbreviation Definition
AD Autonomous Driving
AV Autonomous Vehicle
CAV Connected and Autonomous Vehicle
VIC vehicle-infrastructure cooperation
CP Cooperative or Collaborative Perception
V2X Vehicle-to-Everything
V2V Vehicle-to-Vehicle
V2I Vehicle-to-Infrastructure
V2N Vehicle-to-Network
I2I Infrastructure-to-Infrastructure
RSU Roadside Unit
ITS Intelligent Transportation System
CAM Cooperative Awareness Message
CPM Cooperative Perception Message
CPS Collective Perception Service
CAD Collaborative Autonomous Driving
CDA Cooperative Driving Automation
GPS Global Positioning System
IMU Inertial Measurement Unit
BEV Bird’s Eye View
CNN Convolutional Neural Network
MF Motion Forecasting
Re-ID Re-Identification
LLM Large Language Model
II Collaboration-Based Perception Datasets
With the rapid development of Intelligent Connected Vehicles (ICVs) and Intelligent Transportation Systems (ITS), CP paradigms have become increasingly diverse. These paradigms now encompass not only pure roadside perception but also V2V, V2I, and more comprehensive V2X. Additionally, emerging paradigms such as I2I collaboration have recently gained attention. As illustrated in Fig. 4, various CP architectures are now being explored. In this section, we classify and discuss CP datasets according to their communication paradigms, including roadside perception datasets (Section II-A), V2V datasets (Section II-B), V2I datasets (Section II-C), V2X datasets (Section II-D), and I2I datasets (Section II-E). Table II provides a comprehensive summary of representative datasets under each paradigm. These datasets play a crucial role in advancing cooperative intelligence for ICVs and ITS.
Table II: Publicly Available Datasets for Collaborative Perception
View Dataset Year Venue Sensors Source Tasks Size Link
Image LiDAR 3D Box
Infra Ko-PER [43] 2014 ITSC C, L Real 3DOD, MOT 18.7k 4.8k - Link
CityFlow [44] 2019 CVPR C Real MTSCT/MTMCT, ReID 118k - 22.9k Link
INTERACTION [45] 2019 IROS C, L Real 2DOD, TP 1.4M - 1.4M Link
A9-Dataset [73] 2022 IV C, L Real 3DOD 1.1k - 14.4k Link
IPS300+ [74] 2022 ICRA C, L Real 2DOD, 3DOD 56.7k 14.2k 4.5M Link
Rope3D [75] 2022 CVPR C, L Real 2DOD, 3DOD 50k - 1.5M Link
LUMPI [76] 2022 IV C, L Real 3DOD 200k 90k - Link
TUMTraf-I [77] 2023 ITSC C, L Real 3DOD 4.8k 4.8k 57.4k Link
RoScenes [78] 2024 ECCV C Real 3DOD 1.3M - 21.13M Link
H-V2X [52] 2024 ECCV C, R Real BEV Det, MOT, TP 1.94M - - Link
V2V COMAP [79] 2021 ISPRS L, C Sim 3DOD, SS 8.6k - 226.9k Link
CODD [49] 2021 RA-L L Sim Registration - 13.5k 204k Link
OPV2V [40] 2022 ICRA C, L, R Sim 3DOD 44k 11.4k 232.9k Link
OPV2V+ [80] 2023 CVPR C, L, R Sim 3DOD 11.4k+ - 232.9k+ Link
V2V4Real [53] 2023 CVPR L, C Real 3DOD, MOT, S2R 40k 20k 240k Link
LUCOOP [81] 2023 IV L Real 3DOD - 54k 7k Link
MARS [82] 2024 CVPR L, C Real VPR, NR 1.4M - - Link
OPV2V-H [83] 2024 ICLR C, L, R Sim 3DOD 79k - 232.9k+ Link
V2V-QA [84] 2025 arXiv L, C Real 3DOD, PQA - 18k - Link
CP-UAV [35] 2022 NIPS L, C Sim 3DOD 131.9K - 1.94M Link
V2I CoopInf [48] 2020 TITS L, C Sim 3DOD 10k - 121.2K Link
DAIR-V2X-C [39] 2022 CVPR L, C Real 3DOD 71k 71k 1.2M Link
V2X-Seq [42] 2023 CVPR L, C Real 3DOT, TP 71k 15k 464k Link
HoloVIC [85] 2024 CVPR L, C Real 3DOD, MOT 100k - 11.47M Link
OTVIC [86] 2024 IROS L, C Real 3DOD 15k - 24.4k Link
DAIR-V2XReid [55] 2024 TITS L, C Real 3DOD, Re-ID 7.3k - - Link
TUMTraf V2X [87] 2024 CVPR L, C Real 3DOD, MOT 5k 2k 30k Link
V2X-Radar [57] 2024 arxiv L, C, R Real 3DOD 40k 20k 350k Link
V2X V2X-Sim [41] 2022 RA-L L, C Sim 3DOD, MOT, SS 60k 10K 26.6K Link
V2XSet [34] 2022 ECCV L, C Sim 3DOD 44k 11.4k 233k Link
DOLPHINS [88] 2022 ACCV L, C Sim 2DOD, 3DOD 42.3k 42.3k 292.5k Link
DeepAccident [51] 2024 AAAI L, C Sim 3DOD, MOT, SS, TP - 57k 285k Link
V2X-Real [54] 2024 ECCV L, C Real 3DOD 171k 33k 1.2M Link
Multi-V2X [89] 2024 arxiv L, C Sim 3DOD, MOT 549k 146k 4.2M Link
Adver-City [50] 2024 arxiv L, C Sim 3DOD, MOT, SS 24k 24k 890k Link
DAIR-V2X-Traj [56] 2024 NIPS L, C Real MF 808k - 1.4M Link
WHALES [90] 2024 arxiv L, C Sim 3DOD 70k 17k 2.01M Link
V2X-R [58] 2024 arxiv L, C, R Sim 3DOD 150.9k 37.7k 170.8k Link
V2XPnP-Seq [91] 2024 arxiv L, C Real Perception and Prediction 208k 40k 1.45M Link
Mixed Signals [92] 2025 arxiv L Real 3DOD - 45.1k 240.6k Link
SCOPE [93] 2024 arxiv C, L Sim 2DOD, 3DOD, SS, S2R 17k 17k 575k Link
I2I Rcooper [60] 2024 CVPR C, L Real 3DOD, MOT 50k 30k 30k Link
InScope [59] 2024 arxiv L Real 3DOD, MOT - 21.3k 188k Link
Note: Sensors: Camera (C), LiDAR (L), Radar (R). Source: Real = collected in the real world; Sim = generated via simulation. Tasks: 2DOD = 2D Object Detection, 3DOD = 3D Object Detection, MOT = Multi-Object Tracking, MTSCT = Multi-target Single-camera Tracking, MTMCT = Multi-target Multi-camera Tracking, SS = Semantic Segmentation, TP = Trajectory Prediction, VPR = Visual Place Recognition, NR = Neural Reconstruction, Re-ID = Re-Identification, S2R = Sim2Real, MF = Motion Forecasting, PQA = Planning Q&A.
Figure 4: Schematic diagram of different collaborative perception methods, including roadside perception, V2V, V2I, V2X, and I2I.
II-A Roadside Datasets
Roadside perception plays a vital role in AD and ITS by offering a global, occlusion-resilient view of traffic participants [94][95][96]. Unlike vehicle-mounted systems, infrastructure-based sensing units—typically installed on poles, gantries, or overpasses—enable the monitoring of complex intersections, dense traffic flows, and long-range environments with enhanced spatial stability and temporal continuity. By supplementing on-vehicle sensing with infrastructure-level observations, roadside perception significantly improves perception robustness in scenarios involving occlusions, blind spots, or congested conditions.
A variety of representative roadside datasets have been introduced to support infrastructure-based perception research. Early datasets such as Ko-PER [43] and CityFlow [44] primarily target intersection environments and person or vehicle Re-ID tasks. INTERACTION [45] expands the behavioral diversity of roadside scenarios by incorporating multi-national data with adversarial and cooperative driving maneuvers. Rope3D [75] provides high-resolution 3D data along with dedicated metrics for evaluating roadside 3D object detection. LUMPI [76] supports multi-view analysis using synchronized multi-sensor setups, while IPS300+ [74] offers dense, long-range 3D annotations for complex urban intersections. TUMTraf-I [77] enhances diversity through coverage of varying lighting and weather conditions. For highway-oriented research, A9-Dataset [73] offers real-world, high-speed traffic scenarios, whereas RoScenes [78] features the largest perception area to date through high-density, multi-camera roadside deployment. H-V2X [52] extends this trend by integrating radar and camera data, supporting a range of tasks including detection, tracking, and trajectory prediction under complex highway conditions.
The release of roadside datasets has significantly broadened the scope of perception research, enabling detailed studies on object detection, semantic segmentation, and multi-object tracking from an infrastructure-centric perspective. These datasets also provide a valuable foundation for future V2I CP systems by enhancing scene coverage and perceptual redundancy. However, several challenges remain inherent to infrastructure-based data collection. Such as, high deployment and maintenance costs restrict large-scale installations, particularly in highway and rural areas. Data distributions are often skewed toward urban environments, which limits the generalizability of trained models across diverse traffic scenarios. Additionally, precise temporal synchronization and spatial alignment among heterogeneous sensors remain non-trivial, especially in long-term or multi-site deployments.
II-B V2V Datasets
V2V CP enhances perception range and accuracy by enabling information sharing among multiple vehicles. Compared to single-vehicle perception, V2V collaboration effectively mitigates blind spots caused by occlusions and compensates for the limitations of individual sensors in specific scenarios. This leads to improved robustness in tasks such as object detection, semantic understanding, and behavior prediction. Furthermore, the integration of multi-sensor data and cross-view information enables V2V systems to perform more adaptively in complex and dynamic environments. Therefore, the development of high-quality V2V CP datasets is of great importance for advancing research in multi-vehicle collaborative autonomous driving systems.
In recent years, V2V datasets have continued to evolve in terms of collaborative modeling and data organization mechanisms. Most datasets, including OPV2V [40], MARS [82], and LUCOOP [81], adopt an offline alignment strategy that unifies data from multiple vehicles into a common coordinate system using global positioning or timestamps. This approach simulates idealized synchronized collaboration scenarios. However, it fails to account for real-world issues such as asynchronous sampling, communication latency, and localization errors during actual V2V communication. To address these limitations, HEAL [83] introduces asynchronous perception and temporal alignment mechanisms, while OPV2V-N [97] employs neural modeling to improve perception robustness. COMAP [79] adopts a task-oriented modeling approach from a multi-agent SLAM perspective. In terms of agent configuration, most datasets support CP among 2 to 4 vehicles. MARS [82] enables data collection across multiple time periods, trajectories, and viewpoints using four connected vehicles. OPV2V+ [80] extends viewpoint variation modeling by supporting primary view selection and perception completion. Regarding task coverage, existing datasets support key tasks such as object detection, tracking, Bird’s Eye View (BEV) semantic segmentation, and path planning. V2V-QA [84] further incorporates multi-modal large language models to support intelligent question answering and path reasoning tasks. IRV2V [98] and CODD [49] focus on asynchronous robustness and multi-view point cloud registration, reflecting the expanding scope of V2V data modeling in terms of collaboration granularity and semantic capabilities.
Despite the notable progress made by V2V CP datasets in enabling multi-vehicle fusion research, several distinct challenges remain. For example, most existing datasets rely on offline stitching techniques for data integration and lack modeling of real communication links, transmission latency, and collaborative constraints. As a result, they fail to capture the uncertainty and dynamic characteristics of information exchange in real V2V scenarios, limiting deeper investigations into the essence of cooperative mechanisms. Looking ahead, the development of V2V datasets should focus on realistic modeling of communication processes, the design of dynamic alignment mechanisms for CP, and the construction of efficient multi-vehicle data acquisition and annotation pipelines in complex urban environments.
II-C V2I Datasets
V2I perception datasets aim to enhance AD systems by enabling collaboration between vehicles and roadside infrastructure. These datasets incorporate multimodal data from both vehicle-mounted and infrastructure-based sensors, offering a more comprehensive and accurate understanding of the environment.
Representative V2I datasets exhibit distinct characteristics and collaborative strategies. CoopInf [48] employs eight infrastructure sensors for multi-view 3D object detection. DAIR-V2X-Seq [42] is the first large-scale sequential V2X dataset using data from both infrastructure and vehicles. HoloVIC [85] constructs holographic intersections with diverse layouts to overcome occlusion and blind spots through multisensor fusion. TUMTraf-V2X [87] emphasizes perception under complex lighting conditions and supports 3D detection and tracking. OTVIC [86] targets high-speed and noisy communication scenarios with multimodal, multi-view perception data.
While these datasets significantly contribute to improving V2I CP, challenges remain. Sensor synchronization is technically demanding, and many datasets offer limited scene diversity or weather coverage. Future efforts should prioritize building more diverse, comprehensive V2I datasets to improve perception robustness in complex and dynamic driving environments.
II-D V2X Datasets
CP empowered by V2X technology facilitates data sharing and joint environmental understanding through communication among diverse agents, including vehicles, infrastructure, pedestrians, and other IoT devices. This paradigm significantly improves perception robustness and driving safety in autonomous systems.
From the perspective of agent structure and heterogeneity, recent V2X datasets reveal an increasing trend towards modeling complex interactions among heterogeneous agents with varying sensing capabilities. Early works such as V2X-SIM [41] support basic cooperative tasks, yet remain limited in agent type diversity and environmental complexity. The DAIR-V2X series [39, 55, 56] significantly expands the agent spectrum by introducing bi-agent collaboration between vehicles and RoadSide Units (RSUs). These datasets illustrate asymmetric collaboration in heterogeneous settings, where infrastructure and mobile vehicles possess distinct sensor setups and motion patterns.
To capture multi-agent temporal dynamics, V2XPnP [91] offers sequential data involving multiple cooperative vehicles, facilitating trajectory prediction in temporally correlated environments. DOLPHINS [88] further enhances agent diversity by simulating tri-vehicle collaboration in adverse conditions such as rain and fog, which supports evaluation of perception robustness under extreme scenarios. V2X-Real [54] emphasizes large-scale data collection across multiple vehicles and infrastructures, reflecting real-world V2X deployments with rich agent-level interactions. Adver-City [50] and DeepAccident [51] introduce environmental challenges such as snow, low visibility, and accident-prone zones, where CP across various agents becomes critical. Meanwhile, V2X-Radar [57] and V2X-R [58] expand the sensor modalities by incorporating 4D radar into the V2X sensing framework, highlighting the potential of cross-agent, cross-modal fusion in enhancing perception fidelity.
Despite these advancements, V2X datasets still face challenges related to scalability, diversity, and task integration. Many datasets lack coverage of complex urban intersections or multi-agent scenarios with mixed mobility patterns. Additionally, existing benchmarks often address specific tasks in isolation, limiting their support for evaluating end-to-end CP systems across heterogeneous agents.
II-E I2I Datasets
As V2X CP continues to evolve, I2I collaboration has emerged as a promising paradigm to address the limitations of single-agent roadside systems. While V2I, V2I and V2X datasets focus primarily on perception enhancement via VIC, their perception range is often constrained by occlusion, especially when large vehicles or urban structures obstruct the view. I2I CP addresses this issue by enabling multiple roadside infrastructure nodes to share and fuse data, thereby reducing blind spots and improving the robustness of perception in complex environments such as intersections and road corridors.
Two pioneering datasets exemplify the advancement of I2I CP. The RCooper [60] is the first real-world, large-scale roadside CP dataset. It includes two typical scenarios—intersection and corridor—with infrastructure agents equipped with multi-type LiDARs and cameras. RCooper provides high-quality annotations, demonstrating the effectiveness of cross-infrastructure cooperation. In contrast, the InScope dataset [59] specifically targets occlusion-aware perception by deploying LiDARs at multiple viewpoints. Additionally, InScope proposes a novel quantitative metric to evaluate anti-occlusion capabilities, providing a rigorous framework for I2I perception evaluation. V2XPnP [91] further expands the paradigm by integrating large-scale sequential data across V2X modes, particularly I2I, offering valuable support for complex CP tasks. Compared to other paradigms, I2I collaboration is more effective in mitigating occlusion in dense traffic environments. However, Current I2I datasets face challenges in generalization, data synchronization, and scene diversity, highlighting the need for scalable deployment strategies and standardized evaluation protocols.
III Simulation vs. Real-World Datasets
Simulation and real-world datasets serve complementary roles in AD research, differing in cost, source, annotation accuracy, realism, and applicability. Simulation datasets, generated through virtual modeling, offer low cost, high scalability, and automatic annotations, enabling the design of diverse scenarios, including rare or hazardous conditions. However, the simulation-to-reality gap limits their generalization capabilities. Real-world datasets, collected from actual environments, provide greater complexity and realism, reflecting sensor noise and dynamic interactions, making them essential for final model validation. Despite their reliability, they are costly and labor-intensive to acquire and annotate, often lacking scenario diversity. As a result, simulation datasets are preferred for algorithm development and pre-training, while real-world datasets remain critical for evaluation and deployment [36].
Table III: Comparison of Simulated and Real Datasets Characteristics
Characteristic Sim Real
Data Collection Cost $\checkmark$ $\times$×
Scene Diversity, Environmental Diversity $\checkmark$ $\times$×
Dynamic Interaction Complexity $\times$× $\checkmark$
Ability to Simulate New Scenes $\checkmark$ $\times$×
Coverage of Extreme/Dangerous Scenes $\checkmark$ $\times$×
Coverage of Long-Tail Scenarios $\checkmark$ $\times$×
Difficulty in Large-Scale Data Collection $\checkmark$ $\times$×
Processing Complexity $\times$× $\checkmark$
Data Labeling Accuracy $\checkmark$ $\times$×
Time Synchronization Precision $\checkmark$ $\times$×
Data Reality and Authenticity $\times$× $\checkmark$
Data Randomness and Noise $\times$× $\checkmark$
High Complexity and Diversity $\times$× $\checkmark$
Repeatability, Algorithm Iteration Speed, High Safety $\checkmark$ $\times$×
High Cost of Updates and Maintenance $\times$× $\checkmark$
Data Privacy Compliance $\checkmark$ $\times$×
Industry Certification and Recognition $\times$× $\checkmark$
III-A Simulated Datasets
III-A1 Simulation Platforms
Simulators are essential for CP research in AD, offering cost-effective, safe, and annotation-efficient alternatives to real-world data collection. They enable accurate generation of object attributes and automated annotations. Open-source platforms such as CARLA [99], SUMO [100], and AirSim [101] are widely used for data synthesis. CARLA supports diverse driving scenarios and multimodal sensors, including RGB, depth, radar, and LiDAR, but its LiDAR simulation has limitations in point cloud intensity and echo dropout modeling. SUMO provides vehicle and pedestrian trajectory simulation and can be co-simulated with CARLA to improve scene fidelity. The OpenCDA [102] framework integrates both platforms, offering standardized scenario management and benchmarks for perception and planning tasks.
III-A2 Simulation Datasets
Simulation datasets serve as vital testbeds for CP research, offering controllable environments, automated annotations, and support for diverse or rare scenarios. Representative datasets are built on CARLA, SUMO, or LiDARsim platforms. COMAP [79] and OPV2V [40] leverage CARLA-SUMO co-simulation for multi-agent perception across intersections and urban scenes, supporting detection, segmentation, and trajectory prediction. Extensions such as OPV2V+ [80] and OPV2V-H [83] improve agent diversity and simulate asynchronous sensing. DOLPHINS [88] and Adver-City [50] incorporate adverse weather to evaluate sensor robustness. V2X-Sim [41], Multi-V2X [89], and DeepAccident [51] expand support for high-resolution sensors, multi-agent scheduling, and accident prediction. These datasets provide critical foundations for benchmarking CP algorithms and exploring Sim2Real transfer, communication modeling, and generalization across scenarios.
III-B Real-World Datasets
Real-world CP datasets provide authentic sensor data and traffic interactions essential for AD research and model validation. DAIR-V2X [39] is the first large-scale multimodal real-world dataset focusing on V2I cooperation, while V2X-Seq [42] extends it with long-sequence data and traffic signal information. V2V4Real [53] is the first large-scale real-world multimodal V2V dataset, offering valuable support for Sim2Real studies despite limited CAV diversity. LUCOOP [81], collected with three CAVs using heterogeneous sensors, supports urban driving studies, while MARS [82] expands to four CAVs for improved stability and generalization. HoloVIC [85] and OTVIC [86] focus on V2I setups, with OTVIC simulating real-time vehicle-RSU communication to reflect dynamic transmission challenges in infrastructure-assisted perception.
IV Cooperative perception Scenario Datasets
CP datasets can be categorized according to different traffic environments to address the perception needs of AD systems in various scenarios. In this section, we classify them into three types: intersection, urban street, and highway scenarios. Fig. 5 shows the distribution of representative datasets across these categories. This chapter discusses the key datasets in each scenario and their characteristics, providing essential data support for CP research.
Figure 5: A Venn diagram showing the distribution of CP datasets across intersection, urban street, and highway scenarios. The overlapping areas indicate datasets applicable to multiple scenarios.
IV-A Intersection Datasets
Intersections represent one of the most complex traffic environments, where vehicles, pedestrians, and cyclists converge from multiple directions, leading to significant challenges such as occlusions and dynamic interactions. To enhance perception capabilities and environmental understanding, intersection CP datasets typically employ multi-sensor fusion and cross-view information integration techniques. The following datasets are among the most representative in this category.
Ko-PER [43] is one of the earliest datasets employing fixed LiDAR and cameras to monitor intersection dynamics. IPS300+ [74] improves spatial resolution with high-density 3D annotations, while LUMPI [76] introduces multi-view data under varying weather and lighting. TUMTraf-I [77] further targets annotation precision in adverse conditions, such as night and rain. HoloVIC [85] designs diverse holographic intersections to explore sensor placement strategies for CP. Several datasets incorporate intersection scenarios within broader V2X contexts. DAIR-V2X [39] supports V2I perception in urban crossings; DOLPHINS [88] includes intersection and T-junctions with weather variations; DeepAccident [51] focuses on accident prediction at intersections. V2XPnP [91] offers 163 intersection scenarios, supporting trajectory prediction and object tracking. These intersection datasets collectively emphasize high-precision perception in dense, occlusion-prone environments through multi-view sensing and cross-agent fusion. However, challenges such as dynamic traffic variability, sensor blind zones remain. These intersection datasets collectively emphasize high-precision perception in dense, occlusion-prone environments through multi-view sensing and cross-agent fusion.
IV-B Urban Street Datasets
Urban street CP datasets span V2V, V2I, and V2X paradigms, supporting tasks such as object detection, tracking, and trajectory prediction. CityFlow [44] offers multi-camera tracking data across arterials and residential areas. LUCOOP [81] collects time-synchronized multimodal data from three Connected Autonomous Vehicles (CAVs), including 3D point clouds, city models, and UWB ranging data. MARS [82] further extends to four CAVs under diverse weather, enhancing perception stability and agent collaboration. Several datasets strengthen roadside sensing capabilities. RCooper [60] provides the first large-scale real-world roadside dataset for multi-sensor infrastructure perception. RoScenes [78] delivers dense multi-view 3D data in complex urban traffic, supporting BEV perception. V2XPnP [91] captures diverse urban road scenes with multi-agent trajectories and detailed interactions. Together, these datasets address complex urban traffic environments, offering comprehensive data support for perception, fusion, and prediction in multi-agent V2X systems.
IV-C Highway Datasets
With the growing deployment of ICVs on highways, CP datasets for highway scenarios have become increasingly important for traffic management and AD safety. Highway environments demand long-range detection, lane-level perception, and robust environmental sensing, especially under high-speed conditions where precise synchronization is critical.
Although limited in number, highway CP datasets are gradually expanding. TUMTraf A9 [77] captures multi-weather data from Germany’s A9 highway for 3D object detection. OTVIC [86] introduces the first online-transmission, multi-view, multimodal dataset for V2I cooperative detection at 70–110 km/h. V2V4Real [53] includes 410 km of driving data, with 80% collected in highway environments. RoScenes [78] offers 14 ring-road scenarios with up to 800-meter perception range, enabling blind-spot-free coverage. H-V2X [52] combines radar and cameras for robust BEV perception in adverse weather, marking the first real-world, large-scale highway dataset of its kind. Other datasets such as DAIR-V2X [39], DOLPHINS [88], and CityFlow [44] also include partial highway scenarios, contributing to this domain.
V Datasets by Sensor Modality
Sensor modality diversity plays a pivotal role in enhancing the robustness of CP systems, particularly in complex and dynamic environments [103][104]. As illustrated in Fig. 6, typical sensors employed in CP include cameras, LiDAR, and Radar, each corresponding to distinct perception modalities such as RGB images, dense 3D point clouds, and sparse Radar point clouds. The respective strengths and weaknesses of these sensors are summarized in Table IV, providing a comprehensive overview of their capabilities under varying operational conditions. This section categorizes existing CP datasets based on their acquisition modalities, including image-only, LiDAR-only, and multimodal datasets.
Figure 6: Representative sensor types and corresponding perception modalities. (a) Camera and (b) RGB image; (c) LiDAR and (d) LiDAR point cloud; (e) 4D Radar and (f) sparse Radar point cloud. All data samples are derived from the V2X-Radar [57] dataset.
Image-only datasets, while cost-effective and semantically rich, are limited by poor depth estimation. LiDAR-only datasets offer precise 3D localization but are constrained by high sensor cost and limited semantic information. Multimodal datasets, especially those combining camera and LiDAR, dominate recent research due to their complementary strengths. Moreover, Radar-integrated datasets have gained attention for their resilience to adverse weather, marking a growing trend in sensor fusion for reliable perception [57].
Table IV: The Main Advantages and Disadvantages of Cameras, LiDAR, and Radar
Sensor Advantages Disadvantages
Camera • High resolution, capable of detecting color and texture information, good object classification ability.
• Low cost, mature technology.
• Recognizes traffic signs and lane markings.
• Relies on lighting conditions, reduced performance at night.
• Poor distance and speed measurement capability.
• Affected by weather conditions, performance degrades in adverse weather.
LiDAR • High-precision distance measurement, enables 3D environmental modeling.
• Not affected by lighting conditions, performs well at night.
• Wide field of view detection.
• Typically expensive.
• Performance degrades in rain, snow, and is affected by smoke, dust, and other weather conditions.
• Large data volume, complex processing.
Radar • Unaffected by weather and nighttime conditions, operates in all-weather environments.
• Accurate speed measurement, long detection range.
• Moderate cost, lower than LiDAR.
• Strong penetration (fog, rain, dust), strong anti-interference capability.
• Low resolution.
• Less sensitive to static objects, prone to false detections on metallic objects.
• Sparse detection points, weak reflection from pedestrians, making detection difficult.
• Cannot capture color or image information.
V-A Image-only Datasets
Image-only datasets offer high-resolution visual data, supporting tasks such as object detection and semantic segmentation. However, the absence of depth information limits their performance in occlusion handling, distance estimation, and robustness under varying lighting. CityFlow [44] spans large-scale urban scenarios, supporting cross-camera tracking and vehicle ReID, with extensive spatial and camera coverage. WIBAM [105], captured from infrastructure-mounted cameras, focuses on 2D/3D object detection. RoScenes [78], the largest multi-view roadside dataset to date, targets vision-based BEV perception, addressing challenges in multi-camera fusion and dense traffic scenes. Despite their effectiveness in visual tasks, image-only datasets have limited utility in high-precision localization and CP.
V-B LiDAR-only Datasets
LiDAR-only datasets provide accurate spatial geometry and are widely used for 3D detection, tracking, and map construction. While superior in depth perception, the absence of semantic cues limits their effectiveness in classification. V2V-Sim [47] and CODD [49] offer synthetic LiDAR data for perception and registration tasks. LUCOOP [81], a real-world dataset, supports multi-agent SLAM and urban navigation. Mixed Signals [92] enhances roadside cooperation through diverse LiDAR sensors. InScope [59] improves spatial awareness via distributed LiDAR deployments. Although point cloud datasets excel in 3D localization, their semantic limitations drive the shift toward multimodal fusion datasets in CP research.
V-C Multi-modal Datasets
Multi-modal datasets, particularly those fusing LiDAR and cameras, are central to CP research. By combining LiDAR’s precise 3D spatial sensing with camera-derived semantic richness, these datasets improve environmental understanding and robustness [106][107]. Recently, the integration of radar has further enhanced perception under adverse weather and for long-range detection. This section summarizes key LiDAR-Camera and LiDAR-Radar-Camera datasets.
LiDAR-Camera datasets include both infrastructure-based and cooperative datasets. Fixed-infrastructure datasets such as Ko-PER [43], IPS300+ [74], Rope3D [75], A9-Dataset [73], and LUMPI [76] use roadside LiDAR and cameras to produce accurate 3D annotations. H-V2X [52], RCooper [60], and TUMTraf-I [77] extend perception to challenging conditions such as rain and night. UAV and traffic camera datasets like INTERACTION [45] and BAAI-VANJEE [108] support interaction modeling and driving behavior analysis. CP datasets such as OPV2V [40], IRV2V [98], V2V4Real [53], LUCOOP [81], and MARS [82] collect multi-agent LiDAR-Camera data in both real and simulated settings. OPV2V-N [97] enhances camera robustness, while V2V-QA [84] introduces LLM-based collaborative tasks. CoPerception-UAV [35] supplements V2V data with aerial views. In V2X contexts, DAIR-V2X-C [39], V2X-Seq [42], HoloVIC [85], and OTVIC [86] support multi-modal detection, tracking, and planning. Simulated datasets such as V2X-Sim 2.0 [41], V2XSet [34], DOLPHINS [88], DeepAccident [51], V2X-Real [54], and Multi-V2X [89] further address diverse tasks from accident prediction to V2X data augmentation. LiDAR-Radar-Camera datasets such as T & J [46], OPV2V+ [80], OPV2V-H [83], and V2X-Radar [57] enhance CP under adverse conditions by leveraging radar’s robustness to weather and long-range sensing. In summary, LiDAR-Camera datasets remain the cornerstone of CP research, while LiDAR-Radar-Camera fusion introduces new capabilities for perception in challenging environments.
VI Perception Tasks and Metrics
With the development of CP, various datasets have been introduced to support key perception tasks [109][110] . This chapter categorizes them by task type, including object detection, semantic segmentation, object tracking, and trajectory prediction. Object detection focuses on 3D localization of targets, while semantic segmentation enables pixel-level scene understanding. Object tracking and trajectory prediction involve continuous monitoring and future motion inference of dynamic agents. For each task, we present representative datasets and their corresponding evaluation metrics to guide researchers in selecting appropriate datasets and assessment criteria.
VI-A Object Detection
Object detection is one of the most fundamental tasks in CP, encompassing both 2D and 3D object detection. In V2X scenarios, 3D object detection is particularly critical, often relying on LiDAR point clouds, camera images, or multi-modal data to achieve accurate detection of vehicles, pedestrians, and other objects in dynamic environments.
VI-A1 Datasets for Object Detection
Multiple datasets support V2X object detection across different collaboration modes and sensing modalities. DAIR-V2X-C [39] provides a real-world V2I benchmark for multi-modal detection. V2X-Sim 2.0 [41], offers 3D detection data for cooperative agents. HoloVIC [85] constructs holographic intersections for multi-view 3D sensing. V2X-Real [54] supports both V2I and V2V detection at scale, while V2V4Real [53] focuses on diverse road types for collaborative 3D detection. OPV2V [40] and OPV2V+ [80] simulate multi-agent cooperation, with the latter offering improved detection benchmarks. For roadside perception, RCooper [60] enhances infrastructure-based 3D detection under occlusions using multiple sensor layouts. Rope3D [75], while not a V2X dataset, provides monocular 3D detection benchmarks that benefit cooperative vision research. Overall, these datasets span real and simulated settings, support V2I and V2V collaboration, and offer diverse modalities, forming a solid foundation for V2X object detection studies.
VI-A2 Metrics for Object Detection
3D object detection in CP is primarily evaluated using Average Precision (AP)-based metrics and task-related performance indicators [111][112]. AP quantifies detection accuracy by computing precision at different Intersection-over-Union (IoU) thresholds [113]. Commonly used metrics include AP3D (based on 3D IoU) and APBEV (based on BEV IoU), as introduced by the KITTI benchmark [37].
Most CP datasets adopt AP or mAP as core metrics. RCooper [60] reports AP at multiple IoU thresholds. OPV2V [40] and OPV2V+ [80] evaluate vehicle detection at IoU@0.5 and IoU@0.7, with performance also measured within communication ranges. V2X-Real [54] employs both AP and mAP, using lower IoU thresholds (0.3/0.5) to account for varying object sizes. V2X-Sim 2.0 [41] focuses on BEV-based AP at multiple thresholds. Some datasets integrate communication cost into evaluation. V2V4Real [53] reports AP@0.5 and AP@0.7, and introduces Average MegaByte (AM) to quantify transmission overhead under synchronous and asynchronous conditions. DAIR-V2X [39] incorporates communication efficiency and limits evaluation to objects within the ego vehicle’s ROI. Rope3D [75] enhances AP-based evaluation with additional metrics such as Average Ground Center Similarity (ACS), Average Orientation Similarity (AOS), Average Area Similarity (AAS), and Average Four Ground Points Distance (AGD) to assess localization, orientation, and shape accuracy.
In addition to metric design, a range of benchmark methods have been evaluated across typical CP datasets. Table V summarizes the performance of representative algorithms on DAIR-V2X, V2V4Real, OPV2V, and V2XSet, offering insights into algorithm advancements and dataset challenges.
In summary, AP and IoU remain dominant evaluation standards in CP datasets, while extended metrics—such as BEV AP, communication cost, and asynchronous performance—offer deeper insights into real-world constraints. However, AP treats all missed detections equally, regardless of safety relevance. Future research should explore safety-aware evaluation frameworks tailored to CP tasks.
Table V: Performance Comparison of Representative Methods on Benchmark CP Datasets. All values are reported as AP@0.5 / AP@0.7. “-” indicates results not reported in the corresponding paper.
Method Year PUB Real-world Datasets Simulation Datasets
DAIR-V2X AP@0.5/0.7 Ref. V2V4Real AP@0.5/0.7 Ref. OPV2V AP@0.5/0.7 Ref. V2XSet AP@0.5/0.7 Ref.
F-Cooper [46] 2019 SEC 0.7370/0.5600 [96] 0.6930/0.4320 [96] 0.9070/0.8220 [97] 0.8750/0.7390 [97]
V2VNet [47] 2020 ECCV 0.7097/0.4763 [98] 0.6470/0.3360 [96] 0.8970/0.7400 [96] 0.8710/0.6460 [99]
V2VNetrobust [47] 2021 CoRL 0.6610/0.4860 [96] 0.5500/0.3090 [96] 0.9420/0.8540 [96] - -
V2X-Vit [34] 2022 ECCV 0.7398/0.6150 [102] 0.6590/0.4260 [104] 0.9460/0.8560 [103] 0.9100/0.8030 [99]
AttFuse [40] 2022 ICRA 0.7330/0.5530 [102] 0.7010/0.4540 [104] 0.9050/0.8160 [97] 0.8470/0.7400 [97]
Where2comm [35] 2022 NIPS 0.7520/0.5880 [96] 0.7040/0.4690 [96] 0.9440/0.8550 [96] 0.9130/0.8530 [100]
CoBEVT [114] 2022 CoRL 0.6390/0.5167 [102] 0.6480/0.4040 [104] 0.9330/0.8230 [104] 0.9033/0.8269 [102]
DiscoNet [115] 2021 NIPS 0.7370/0.5840 [96] 0.7360/0.4660 [96] 0.9160/0.7910 [103] 0.9078/0.8381 [102]
FPV-RCNN [116] 2022 RAL 0.6550/0.5050 [103] 0.7010/0.4790 [96] 0.8580/0.8400 [103] 0.8650/0.5630 [99]
CoAlign [117] 2023 ICRA 0.7460/0.6040 [103] 0.7090/0.4170 [96] 0.9660/0.9120 [103] 0.9190/0.8050 [96]
SCOPE [93] 2023 ICCV 0.6518/0.4989 [105] - - 0.8752/0.7505 [105] 0.8971/0.8062 [105]
V2VFormer [103] 2024 TIV - - 0.6540/0.3610 [107] 0.9170/0.8760 [107] - -
DI-V2X [118] 2024 AAAI 0.7882/0.6616 [98] - - - - 0.9270/0.8270 [98]
V2X-PC [119] 2024 arXiv 0.7689/0.6939 [101] - - - - 0.9283/0.8955 [101]
VI-B Semantic Segmentation
Semantic segmentation is a crucial task in CP, aiming to achieve pixel-level classification for enhanced environmental understanding. CP semantic segmentation datasets typically integrate LiDAR point clouds and camera images to enable finer scene interpretation, thereby improving AD systems’ perception in complex environments.
VI-B1 Datasets for Semantic Segmentation
Several datasets support semantic segmentation in CP tasks. V2X-Sim 2.0 [41] provides multi-agent LiDAR segmentation with dynamic obstacle, lane, and BEV semantic annotations. OPV2V [40] supports BEV semantic segmentation using LiDAR-camera fusion, while OPV2V+ [80] enhances cross-agent temporal alignment. OPV2V-N [97] focuses on improving image segmentation robustness. COMAP [79] offers BEV semantic segmentation in simulated urban scenarios. DeepAccident [51] provides accident scene segmentation with fine-grained semantic labels. DOLPHINS [88] supports V2X BEV segmentation under varied weather. Adver-City [50] collects segmentation data under extreme weather using dedicated semantic cameras.
Despite these advances, semantic segmentation in CP still faces several challenges, including inconsistent semantic representations across agents, and reduced robustness in adverse weather or sparse data conditions. Moreover, current datasets are limited in diversity and scale, especially for real-world multi-agent segmentation benchmarks. Future trends will emphasize the construction of large-scale real-world V2X segmentation datasets, cross-agent temporal consistency, and robustness under extreme conditions to support reliable and generalizable perception in complex collaborative scenarios.
VI-B2 Metrics for Semantic Segmentation
Semantic segmentation in CP tasks is commonly evaluated using Pixel Accuracy (PA), mean Pixel Accuracy (mPA), Intersection over Union (IoU), and mean IoU (mIoU) [120, 121]. PA measures the ratio of correctly classified pixels, mPA averages PA across classes to address imbalance, and mIoU reflects overall segmentation accuracy by averaging per-class IoU.
In CP scenarios, evaluation extends beyond single-agent performance to assess cross-agent consistency and modality alignment. V2X-Sim 2.0 [41] evaluates BEV semantic segmentation across categories such as lanes, vehicles, and buildings using mIoU. OPV2V-N [97] focuses on map-level segmentation (drivable area, lanes, vehicles), using IoU as the core metric. Effective CP segmentation evaluation requires traditional metrics and additional measures such as temporal stability, cross-modal IoU, and multi-agent coherence.
VI-C Tracking and Prediction
VI-C1 Datasets for Tracking and Prediction
Object tracking and trajectory prediction focus on modeling the temporal dynamics of moving agents and are essential for AD. CP improves tracking accuracy and trajectory prediction by enabling inter-agent information sharing, particularly under occlusion and complex traffic conditions.
Representative datasets include V2X-Seq [42], V2XPnP-Seq [91], and DAIR-V2X-Traj [56]. V2X-Seq offers long-sequence data for both V2V and V2I scenarios, while V2XPnP-Seq enhances temporal consistency and multi-view target association. DAIR-V2X-Traj focuses on V2I trajectory prediction with high-precision annotations. DeepAccident [51] addresses accident prediction by providing trajectory data in critical traffic scenarios.In urban-scale environments, CityFlow [44] supports cross-camera object tracking for traffic monitoring, and INTERACTION [45] provides detailed vehicle trajectories for interactive behavior modeling. RCooper [60] employs roadside sensors to support tracking in dense traffic. OPV2V [40] and V2V4Real [53] benchmark V2V tracking and trajectory prediction in simulated and real-world settings, respectively. V2X-Sim 2.0 [41] supports multi-agent tracking and prediction using BEV-based annotations.
Overall, CP datasets for tracking and prediction span V2V, V2I, and roadside sensing paradigms, providing diverse benchmarks for multi-agent temporal perception and forecasting in AD systems.
VI-C2 Metrics for Tracking and Prediction
Object tracking and trajectory prediction are central to CP, requiring precise evaluation of detection accuracy, tracking stability, and motion forecasting. Standard metrics have been developed to meet these requirements, with extensions introduced to address the complexities of multi-agent V2X scenarios.
For tracking tasks, the CLEAR MOT framework is widely adopted, incorporating Multiple Object Tracking Accuracy (MOTA), Precision (MOTP), and Identity F1 Score (IDF1). MOTA considers False Positives (FP), False Negatives (FN), and Identity Switches (IDSW), while MOTP measures average IoU between predicted and ground-truth trajectories. IDF1 evaluates consistency in identity association. Several datasets extend this framework. RCooper [60] and V2V4Real [53] incorporate Average MOTA (AMOTA), Average MOTP (AMOTP), and scaled AMOTA (sAMOTA) to assess tracking under varied object sizes and detection thresholds. DAIR-V2X-Traj [56] and V2X-Seq [42] also report MOTA, MOTP, and ID Switches for comprehensive analysis. To better reflect multi-agent collaboration, V2X-Sim 2.0 [41] introduces Higher Order Tracking Accuracy (HOTA), which jointly evaluates detection, association, and temporal stability. It also uses Byte Per Second (BPS) to quantify transmission cost, enabling analysis of the trade-off between bandwidth and accuracy.
Trajectory prediction tasks primarily use Average Displacement Error (ADE) and Final Displacement Error (FDE). ADE measures the mean offset over the entire predicted trajectory, while FDE focuses on the final point prediction accuracy. Datasets like INTERACTION [45] and V2X-Seq [42] adopt these metrics, with V2X-Seq further introducing Missing Rate (MR) to capture trajectory coverage completeness. In accident prediction, DeepAccident [51] proposes Volumetric Prediction Quality (VPQ) to jointly evaluate object detection and trajectory forecasting. It also introduces Accident Prediction Accuracy (APA), based on True/False Positives and Negatives, offering a safety-aware evaluation scheme for high-risk scenarios.
Overall, MOTA, MOTP, ADE, and FDE remain core metrics for tracking and prediction, while newer metrics like HOTA, BPS, VPQ, and APA enable more comprehensive evaluations tailored to V2X CP. These frameworks offer critical tools for benchmarking CP performance and guiding future system optimization.
VII Comprehensive Analysis of Collaborative Perception Datasets
To support robust and scalable development of CP systems, it is essential to comprehensively understand the characteristics, accessibility, and research impact of existing datasets [122]. This section systematically analyzes all reviewed CP datasets from three perspectives: their public availability and access mechanisms, the overall quality in terms of modality, annotation, and scene diversity, and the influence of representative datasets on the research community. These insights provide valuable references for dataset selection, algorithm benchmarking, and future dataset construction.
VII-A Availability and Access Methods of Datasets
Dataset accessibility plays a critical role in promoting research progress in CP [123] . Publicly available datasets reduce annotation costs and enable reproducibility, fostering innovation and collaboration. To assess the openness of current datasets, we categorize existing CP datasets based on their access methods.
First, a large portion of datasets can be directly downloaded from official websites, offering high accessibility without requiring registration or application. These include Ko-PER [43], CityFlow [44], LUMPI [76], TUMTraf-I [77], RoScenes [78], OPV2V [40], OPV2V+ [80], V2V4Real [53], LUCOOP [81], TUMTraf V2X [87], V2X-Sim 2.0 [41], DeepAccident [51], V2X-Real [54], Multi-V2X [89], Adver-City [50], MARS [82], Mixed Signals [92], and SCOPE [93]. In addition, some datasets are hosted on GitHub, often accompanied by source code, benchmark tools, or baseline implementations. Representative examples include V2V-Sim [47], OPV2V-H [83], CoPerception-UAV [35], CoopInf [48], V2XSet [34], and V2X-R [58], which facilitate rapid experimentation and method replication. Meanwhile, several datasets are distributed via cloud platforms such as Baidu Netdisk, Google Drive, or Alibaba Cloud. Access typically requires a request or registration process. These include H-V2X [52], DAIR-V2X-C [39], V2X-Seq [42], OTVIC [86], V2XReid [55], V2X-Radar [57], WHALES [90], and V2XPnP-Seq [91]. Furthermore, a number of datasets require formal application procedures, often involving submission of research proposals or institutional endorsements. Examples include BAAI-VANJEE [108], Rope3D [75], A9-Dataset [73], HoloVIC [85], V2V-QA [84], CODD [49], DAIR-V2X-Traj [56], INTERACTION [45], IPS300+ [74], DOLPHINS [88], InScope [59], and RCooper [60]. Finally, a few datasets do not publicly disclose their access procedures, possibly due to ongoing internal usage, limited availability, or institutional restrictions. These include WIBAM [105], T & J [46], IRV2V [98], OPV2V-N [97], CARTI [124], and V2X-DSI [125].
In summary, while many CP datasets are openly accessible via official websites, GitHub, or cloud storage, a significant portion remains gated behind application-based processes or lacks clear documentation. Improving data openness, unifying access procedures, and providing transparent distribution mechanisms will be essential to advancing CP research.
VII-B Comprehensive Quality Analysis of Datasets
Figure 7: Qualitative radar chart analysis of CP datasets. Each subplot represents a category of datasets and compares representative datasets across six key dimensions: Frame (number of frames captured in each dataset), Labels (annotation quality and category diversity), Agents (number and heterogeneity of collaborating entities), Scenarios (scene types and environmental coverage), Modalities (sensor types: camera, LiDAR, radar), and Tasks (supported perception tasks: object detection, tracking, segmentation, prediction). The analysis is qualitative and intended to provide an intuitive comparison of dataset characteristics, highlighting their strengths and focus areas in different CP contexts.
The quality of CP datasets directly impacts their applicability in real-world scenarios. As summarized in Table II and illustrated qualitatively in Fig. 7, we conduct a multi-dimensional analysis across six key aspects: data frame volume, annotation richness, agent diversity, scenario complexity, sensor modalities, and supported tasks. These dimensions reflect both dataset scalability and task coverage, providing a unified view for comparative assessment.
In terms of data volume, large-scale datasets such as RoScenes [78], INTERACTION [45], and DAIR-V2X-Traj [56] contain millions of frames or annotations, supporting long-sequence modeling and prediction tasks. For sensor diversity, OPV2V [40], V2X-R [58], and V2X-Radar [57] integrate camera, LiDAR, and radar modalities, enabling robust multimodal fusion. Regarding task support, datasets like V2X-Sim [41], DeepAccident [51], and SCOPE [93] offer multiple annotations for detection, tracking, segmentation, and prediction, enhancing their versatility for perception research. The radar charts in Fig. 7 further visualize dataset strengths and limitations across these dimensions within categories such as V2V, V2I, and roadside perception.
Overall, high-quality datasets exhibit complementary characteristics: real-world datasets like DAIR-V2X [39] and V2X-Real [54] emphasize fidelity and generalization, while simulation-based datasets like OPV2V+ [80] and Multi-V2X [89] excel in scale, flexibility, and controlled benchmarking. The increasing inclusion of diverse modalities, rich annotations, and comprehensive task coverage highlights a clear trend toward more holistic, multi-agent CP datasets. Future research would benefit from standardized benchmarking protocols across these high-quality datasets to promote reproducibility and cross-domain evaluation.
VII-C Summary of High-Influence Datasets
In the field of CP, several milestone datasets have gained prominence. This section summarizes key high-impact datasets across infrastructure-based, V2I, V2V, and V2X CP domains.
High-impact datasets are typically defined by high citation rates and widespread use in benchmark evaluations. For infrastructure-based perception, CityFlow [44], INTERACTION [45], and Rope3D [75] are widely used in multi-camera tracking and urban scene understanding. Representative methods such as BEVHeight [126], BEVHeight++ [127], and ViT-CoMer [128] have been developed using these datasets. In V2V perception, OPV2V [40] is one of the most cited simulation datasets, supporting algorithms like Where2comm [35], CoBEVT [114], and OPV2V-H [83]. In the V2I domain, DAIR-V2X [39] and its extensions (e.g., DAIR-V2X-Seq [42], DAIR-V2X-Traj [56]) provide large-scale real-world data with high-quality multimodal annotations, supporting tasks such as detection, tracking, and motion forecasting. These datasets underpin studies including CoCa3D [80], HM-ViT [129], BEV-V2X [130], and How2comm [131]. In the V2X category, V2X-Sim [41] supports BEV detection and semantic segmentation, and is frequently adopted by methods such as VoxFormer [132] and V2X-ViT [34].
In summary, datasets like CityFlow, INTERACTION, Rope3D, OPV2V, DAIR-V2X, and V2X-Sim have become cornerstones of CP research due to their broad adoption and benchmark availability. Newer datasets, including V2V4Real [53], Multi-V2X [89], TUMTraf V2X [87], and RCooper [60], are also gaining momentum and are expected to play increasingly important roles.
VIII Challenges and Frontiers
As research on CP continues to advance, datasets play a crucial role in facilitating multi-agent perception, V2X communication, and intelligent transportation systems. However, existing CP datasets still face numerous challenges, including insufficient data scale and diversity, limited availability, lack of standardization, domain adaptation issues, and concerns over privacy and security. Furthermore, with the rapid development of LLMs, their application in AD and CP datasets introduces new research trends. This chapter discusses these core challenges and explores potential future directions.
VIII-A Dataset Scale and Diversity
VIII-A1 Large-Scale Datasets
Large-scale datasets are essential for training deep learning-driven CP systems and improving their generalization ability. However, existing CP datasets are still significantly smaller than large-scale datasets in single-vehicle perception, such as the Waymo Open Dataset and nuScenes. While datasets like OPV2V, DAIR-V2X, and V2X-Sim provide relatively large sample sizes, they remain insufficient to comprehensively cover the complex and dynamic real-world traffic environment. Additionally, constructing large-scale datasets is costly and involves complex annotation processes, limiting their scalability. Future research may leverage automated annotation, synthetic data generation, and data augmentation techniques to reduce construction costs while enhancing dataset diversity and representativeness to better meet the needs of CP tasks.
VIII-A2 Datasets for Adverse Environment
Most current CP datasets primarily cover scenes under clear weather and good lighting conditions, with limited data available for adverse weather scenarios such as heavy rain, snow, fog, and extreme lighting. Adverse environments significantly impact sensor perception capabilities, such as LiDAR experiencing severe signal attenuation in rain and cameras suffering performance degradation under strong or low-light conditions. The lack of such data leads to poor generalization of existing perception models in extreme environments. Future research should focus on collecting data in adverse weather conditions and developing weather simulation techniques in virtual environments to supplement missing data. Additionally, leveraging multimodal sensor fusion and robustness-enhancing algorithms can improve model stability under challenging conditions.
VIII-A3 Datasets for Complex Scenarios
CP must handle complex scenarios such as high-density traffic, intersections, and lane changes. However, current datasets provide limited coverage of these scenarios. Object occlusion, dynamic interactions, and uncertainties in complex environments increase the difficulty of perception and prediction tasks. Future datasets should include diverse road types and driving behaviors to better simulate real-world driving conditions. Enhancing trajectory prediction and multi-agent interaction annotations will also provide a reliable benchmark for evaluating the performance of CP systems in complex scenarios.
VIII-A4 Datasets for Extreme Scenarios
Extreme scenarios are critical for evaluating the safety of AD systems. However, existing datasets contain few extreme scenario samples, limiting model generalization in these key situations. For example, the DeepAccident dataset attempts to provide traffic accident data, but its coverage of accident types remains incomplete. Future research should explore generating extreme scenario data through simulation, combined with real-world data augmentation, to enhance model responsiveness in hazardous situations. Moreover, anomaly detection and Out-Of-Distribution (OOD) analysis can help identify extreme scenarios, improving the safety and stability of CP systems.
VIII-A5 Geographic Diversity in Datasets
Traffic rules, driving behaviors, and road structures vary significantly across different countries and regions, posing challenges for deploying CP models in diverse geographical areas. Many publicly available datasets are region-specific; for instance, DAIR-V2X is primarily based on Chinese road data. The lack of geographic diversity may lead to performance degradation when models are deployed in different countries. Future datasets should incorporate traffic data from various regions to enhance geographic diversity and improve cross-region generalization capabilities. Additionally, transfer learning and domain adaptation techniques can help models better adapt to different traffic environments.
VIII-B Dataset Availability and Accessibility
Although multiple CP datasets are publicly available, some remain restricted, requiring special permissions or being limited to specific research collaborations. Furthermore, dataset storage and distribution methods are inconsistent, with some datasets hosted on unstable platforms, making access difficult for researchers. To promote research fairness and reproducibility, future efforts should focus on establishing more open data-sharing mechanisms, such as cloud-based platforms with unified data access interfaces. Additionally, standardizing dataset storage formats can improve portability and compatibility, allowing different research teams to efficiently utilize available data.
VIII-C Dataset Standardization and Benchmarking
Current datasets exhibit significant differences in annotation formats, category definitions, and evaluation metrics, making direct comparisons between studies challenging. For example, some datasets adopt the KITTI format for 3D object detection annotations, while others follow the nuScenes or Waymo formats. Additionally, inconsistencies in coordinate systems, units, and category definitions hinder data interoperability. Establishing a unified dataset standard that includes annotation specifications, data formats, and evaluation metrics will enhance compatibility and reproducibility, promoting standardized development in CP research.
VIII-D Domain Adaptation and Generalization
Domain adaptation remains a key challenge in improving dataset generalization. Due to differences in data collection environments, sensor configurations, and traffic scenarios, models trained on one dataset may not perform well on another. For instance, CP models trained in simulation environments often struggle when applied to real-world scenarios. Future research should explore unsupervised domain adaptation and self-supervised learning methods to enhance cross-domain generalization. Additionally, using Generative Adversarial Networks (GANs) to synthesize cross-domain data and leveraging large-scale pre-trained models for domain transfer learning could provide effective solutions.
VIII-E Privacy and Security
CP involves multi-agent data sharing, raising concerns about privacy protection and data security. Malicious attackers could intercept vehicle sensor data or manipulate transmitted data to disrupt collaborative decision-making. Furthermore, datasets may contain sensitive information, such as vehicle identification details and pedestrian facial features, increasing privacy risks. To address these issues, techniques such as differential privacy and federated learning can be employed to protect user privacy while maintaining data usability. Additionally, ensuring secure V2X communication mechanisms to prevent cyber attacks is an important direction for future research.
VIII-F LLMs and Datasets
Recently, LLMs have gained attention in the field of AD and intelligent transportation. LLMs can be used for automatic data annotation, synthetic data generation, and traffic scene analysis. However, their integration with V2X datasets is still in the early exploration stage. Key research directions include leveraging LLMs to enhance annotation efficiency and integrating LLMs for semantic understanding in multimodal perception tasks. Future trends may involve LLM-based data augmentation, AD knowledge graph construction, and the fusion of natural language with perception tasks to improve the intelligence level of V2X perception systems.
In summary, current CP datasets still face numerous challenges in terms of scale, availability, standardization, domain adaptation, and privacy security. Future directions include expanding dataset diversity and scale, establishing more comprehensive data-sharing mechanisms, promoting standardization efforts, and exploring the application of emerging technologies such as LLMs in dataset development. Advancements in these areas will further drive CP research and enhance the safety and reliability of AD systems.
IX Conclusion
This review presents a comprehensive review of CP datasets for AD, classifying them by collaboration paradigms, sensor modalities, data collection methods, environmental scenarios, and supported tasks. We analyze both real-world and simulation-based datasets across V2V, V2I, V2X, and I2I settings, evaluating their annotation quality, accessibility, scenario coverage, and benchmarking readiness. While existing datasets have significantly advanced CP research, limitations remain in data scale, coverage of rare and adverse conditions, agent diversity, and geographic generalization. To bridge these gaps, future datasets should incorporate long-tail scenarios such as traffic anomalies and extreme weather, include data from diverse regions, support heterogeneous agent types, and reflect realistic communication constraints. Establishing standardized annotation formats and unified evaluation protocols is essential for enabling fair comparisons and fostering cross-dataset generalization. Additionally, techniques such as automated labeling, high-fidelity simulation, and generative modeling—especially with the aid of large language models—offer promising directions for building scalable and semantically rich datasets. We hope this review provides valuable guidance for the design, construction, and benchmarking of future datasets, ultimately advancing CP systems in real-world AD applications.

### v2

@@
 I Introduction
-Perception serves as a foundational component of Autonomous Driving (AD) systems [1], [2], [3], [4], [5], [6], [7], [8], [9], enabling the vehicle to perceive its surroundings through various onboard sensors and to perform real-time tasks such as object detection [10][11], semantic segmentation, tracking, and trajectory prediction in dynamic traffic environments [12].
-The precision and reliability of the perception system directly impact the safety and decision-making capability of autonomous vehicles, making it a critical aspect to achieve high-level AD.
-In recent years, significant progress has been made in single-vehicle perception, driven by progress in computer vision [13][14][15], deep learning [16][17], and sensor technology [18][19].
-Numerous studies have leveraged cameras [20][21][22], LiDAR [23][24], Radar [25][26], and multi-sensor fusion [27][28][29][30][31][32] to enhance environmental perception in AD systems.
-However, traditional single-vehicle perception is constrained by limited fields of view, sensor blind spots, and environmental occlusions, making it insufficient to meet the high-reliability requirements of AD in complex and dynamic traffic environments.
+Perception is fundamental to Autonomous Driving (AD) systems [1][2][3], [4][5][6][7][8], enabling vehicles to understand their surroundings via onboard sensors for tasks such as object detection [9][10], semantic segmentation, tracking [11], and trajectory prediction [12].
+The accuracy and robustness of perception directly affect safety and decision-making.
+Recent advances in computer vision [13], [14], [15], deep learning [16][17], and sensor technology [18], [19][20] have led to substantial progress in single-vehicle perception, leveraging cameras [21], [22], LiDAR [23][24], radar [25], [26], and sensor fusion [27][28][29], [30], [31].
+However, these systems are constrained by limited field of view, sensor blind spots, and occlusions, making them insufficient in complex, dynamic traffic scenarios.
 Figure 1: Development Trends of Collaborative Perception Datasets.
-To overcome the limitations of single-vehicle perception, Collaborative Perception (CP) leverages Vehicle-to-Everything (V2X) communication to enable real-time information sharing among vehicles and infrastructure [33], thereby enhancing the perception of occluded and long-range regions, improving driving safety, and expanding the Operational Design Domain (ODD) [34][35].
-For instance, at an urban intersection, a single vehicle may fail to detect pedestrians or oncoming vehicles that are occluded by buildings or other obstacles.
-By leveraging V2X connectivity, autonomous vehicles can receive supplementary perception data from roadside infrastructure or neighboring vehicles, thereby improving situational awareness and reducing accident risk.
-V2X communication encompasses Vehicle-to-Vehicle (V2V), Vehicle-to-Infrastructure (V2I), Infrastructure-to-Infrastructure (I2I), and Vehicle-to-Network (V2N) links, empowering AD systems to overcome the inherent limitations of individual perception and achieve a more holistic understanding of their surroundings.
-High-quality datasets are essential for advancing algorithm development and performance evaluation in CP research [36].
-Existing single-vehicle perception datasets, such as KITTI [37], nuScenes [38], and Waymo [2], are primarily designed for individual perception tasks and do not support multi-vehicle or infrastructure-level information exchange, rendering them unsuitable for direct use in CP studies.
-In recent years, numerous CP datasets, as illustrated in Fig.
+To address these limitations, Collaborative Perception (CP) leverages Vehicle-to-Everything (V2X) communication to enable real-time data sharing among vehicles and infrastructure [32], thereby enhancing the perception of occluded and long-range regions, improving safety, and expanding the Operational Design Domain [33], [34].
+For example, at urban intersections, a vehicle may fail to detect occluded agents; V2X links allow perception data exchange with infrastructure or other vehicles to overcome such blind spots.
+V2X includes Vehicle-to-Vehicle (V2V), Vehicle-to-Infrastructure (V2I), Infrastructure-to-Infrastructure (I2I), and Vehicle-to-Network modes, enabling more comprehensive environmental awareness.
+High-quality datasets are critical for algorithm development and benchmarking in CP [35].
+Traditional datasets like KITTI [36], nuScenes [37], OpenMPD [38], and Waymo [2] focus on single-agent perception and lack support for inter-agent communication, making them unsuitable for CP research.
+Recently, numerous CP datasets, as illustrated in Fig.
 1, have been released to support various scenarios, tasks, and sensor modalities, including DAIR-V2X [35], OPV2V [39], V2X-Sim [40], and V2X-Seq [41].
-While these datasets offer valuable resources for the CP community, they differ significantly in terms of scale, scene diversity, annotation quality, sensor configurations, and the simulation of communication latency.
-A systematic review and analysis of existing CP datasets can help researchers better understand their characteristics, application boundaries, and limitations, thereby providing scientific guidance for future investigations.
-Furthermore, identifying existing gaps and prospective research directions will contribute to the continued advancement of CP technologies.
+These datasets vary in scale, annotation quality, sensor layout, and latency modeling.
+A systematic review is necessary to clarify their characteristics, limitations, and application scope, and to inform future research directions in collaborative perception.
 I-A Roadmap
-Figure 2: Chronological overview of collaborative perception datasets, including roadside perception, V2V, V2I, V2X, and I2I datasets.
+Figure 2: Chronological overview of collaborative perception datasets, including roadside, V2V, V2I, V2X, and I2I datasets.
 Fig.
-2 illustrates the chronological evolution of CP datasets.
-Overall, the development of these datasets has progressed through several key phases, each marking a significant shift in perception paradigms or improvements in dataset diversity.
-Changes in collaboration models, increasing scene complexity, the expansion of supported tasks, and the continuous growth in sensor configurations have all contributed to the advancement of CP technologies.
-In the early stages, datasets were primarily focused on infrastructure-based perception, as exemplified by Ko-PER [43], CityFlow [44], and the INTERACTION [45].
-However, these datasets did not support V2V or V2I information exchange, thereby limiting their applicability to CP research.
-With the advancement of V2X communication technologies, the number of related datasets has steadily increased.
-Around 2020, several V2V and V2I datasets were released, including T & J [46], V2V-Sim [47], CoopInf [48], and CODD [49], which collectively accelerated progress in CP research.
-However, datasets from this period still faced limitations in terms of scene complexity, sensor modality diversity, and task coverage.
-As the field progressed, newer CP datasets began to target a broader range of scenarios.
-For example, Adver-City [50] focuses on adverse weather conditions, DeepAccident [51] addresses accident prediction, and H-V2X [52] is tailored for highway environments.
-These datasets have significantly enhanced the adaptability of CP systems.
-Since 2022, CP dataset development has increasingly emphasized multi-modal fusion and real-world deployment.
-Datasets such as V2V4Real [53], V2X-Real [54], and DAIR-V2X-C [39] incorporate real-world data collection to overcome the limitations of synthetic datasets in terms of authenticity and distribution.
-Simultaneously, newer datasets demonstrate broader task coverage and richer sensor modalities.
-For instance, V2X-Seq [42], V2XReID [55], and DAIR-V2X-Traj [56] support a variety of perception tasks, while V2X-Radar [57] and V2X-R [58] incorporate 4D radar to enhance perception quality.
-Moreover, the emergence of I2I datasets such as InScope [59] and Rcooper [60] has further enriched the collaborative paradigms within CP.
-In general, the evolution of CP datasets reflects a continuous trend toward greater complexity, richer modalities, and more specialized scenarios.
+2 summarizes the chronological evolution of CP datasets across major collaboration paradigms.
+Development has progressed from static roadside perception to dynamic multi-agent collaboration, driven by increasing task complexity, richer sensors, and advanced communication technologies.
+Initial datasets, such as Ko-PER [42], CityFlow [43], and INTERACTION [44], focused on roadside perception without inter-agent communication.
+Around 2020, datasets like T&J [45], V2V-Sim [46], CoopInf [47], and CODD [48] introduced V2V and V2I collaboration, though often under constrained settings.
+Later efforts expanded scenario coverage: Adver-City [49] targets adverse weather, DeepAccident [50] supports motion and accident prediction, and H-V2X [51] covers highway scenes.
+Since 2022, the focus has shifted to real-world deployment and sensor diversity.
+Datasets such as V2V4Real [52], V2X-Real [53], and DAIR-V2X-C [35] emphasize realism, while V2X-Seq [41], V2XReID [54], and DAIR-V2X-Traj [55] extend support to trajectory and identity tasks.
+V2X-Radar [56] and V2X-R [57] incorporate 4D radar, and V2XSet-Noise [58] simulates deployment noise for robustness benchmarking.
+Recent I2I datasets such as InScope [59] and RCooper [60] further diversify collaboration modes.
+Overall, CP datasets are evolving toward higher realism, richer modalities, and more specialized tasks.
 I-B Comparison to Related Reviews
-We aim to clarify the distinctions between this review and previous related reviews [61], [36], [62], [63], [64], [65], [66], [67], [68], [69], [70].
-While some existing reviews partially overlap with ours in terms of dataset coverage [71] [72], they omit several important CP datasets and recent benchmark methodologies introduced in the past few years.
-Furthermore, their classification frameworks tend to be narrow and often lack systematic summarization and in-depth comparative analysis.
-In addition, some reviews concentrate primarily on vehicle-centric datasets [61][36][62][63], or explore specific aspects of CP such as multimodal fusion [64][65] or certain CP techniques [66][67][68][69][70], without offering a comprehensive overview of dataset evolution and their applicability across various scenarios.
-In contrast, this review presents the most up-to-date and complete collection of CP datasets, encompassing a broader spectrum of perspectives.
-It also delivers an in-depth discussion on dataset diversity, key challenges, and emerging research directions.
+This review distinguishes itself from prior reviews [61], [62], [63], [64], [65], [66], [67], [68], [69], [70], [71], which primarily focus on vehicle-centric datasets, specific techniques, or particular fusion strategies.
+While a few works partially overlap with ours in dataset coverage [72], [73], they overlook several recent CP datasets and benchmarks, and often lack systematic categorization or comparative depth.
+In contrast, our review provides a comprehensive and up-to-date compilation of CP datasets, with detailed analysis of their diversity, limitations, and suitability across various CP scenarios.
 Figure 3: Organization of this paper.
 I-C Main Contributions
-In this review, we present a systematic and comprehensive analysis of CP datasets for AD, as summarized in Fig. 3. The main contributions of this work are as follows:
-Comprehensive Review of CP Datasets: To the best of our knowledge, this is the first dedicated and comprehensive review focused specifically on CP datasets.
-We systematically categorize and analyze available datasets based on collaboration types, data sources, sensor modalities, and application scenarios.
-Task-Oriented Dataset Classification: We classify CP datasets according to core perception tasks in AD, including 3D object detection, semantic segmentation, and object tracking and prediction.
-This task-centric taxonomy facilitates effective dataset selection based on research needs.
-In addition, we cover the latest developments in CP datasets, providing a timely and up-to-date overview.
-Dataset Comparison and Emerging Trends: We conduct multi-dimensional comparisons, highlight high-impact datasets, and discuss key challenges and future directions such as standardization, LLM-driven annotation, and data diversity.
+The key contributions are summarized as follows:
+Comprehensive Dataset Review: To the best of our knowledge, this is the first dedicated review that focuses exclusively on CP datasets.
+We systematically categorize and comprehensively analyze the existing datasets based on collaboration paradigms, data sources, sensor modalities, and application scenarios.
+Task-Centric Classification: We organize datasets by core perception tasks in AD, including 3D object detection, semantic segmentation, and object tracking and prediction.
+This task-oriented taxonomy aids researchers in selecting appropriate datasets for specific applications.
+Recent developments and newly released datasets are also included to ensure timeliness.
+Comparative Analysis and Future Trends: We conduct multi-dimensional comparisons, highlight representative datasets, and discuss emerging challenges such as standardization, LLM-based annotation, and cross-domain diversity, providing insights for future CP research.
 I-D Review Structure
-The structure of this paper is illustrated in Fig.
+The organization of this paper is illustrated in Fig.
 3.
-To better present our review, the remainder of this paper is organized as follows.
-Section II introduces datasets under different collaborative paradigms.
-Section III compares the characteristics of simulated and real-world datasets.
-Section IV reviews existing CP datasets from the perspective of different application scenarios.
-Section V analyzes datasets based on sensor modalities.
-Section VI summarizes and analyzes the CP tasks supported by these datasets, along with their corresponding evaluation metrics.
-Section VII provides a comprehensive analysis of existing CP datasets, focusing on their availability, quality, and impact.
-Section VIII discusses the challenges and trends in CP datasets.
-Section IX concludes the paper and outlines future research directions.
-A summary of all abbreviations used throughout the paper is provided in Table I.
+Section II introduces CP datasets grouped by collaboration paradigms.
+Section III compares simulated and real-world datasets.
+Section IV focuses on application scenarios, while Section V analyzes sensor modalities.
+Section VI covers core perception tasks and evaluation metrics.
+Section VII provides dataset-level comparisons.
+Section VIII discusses key challenges and trends.
+Section IX concludes the paper.
+A summary of abbreviations is provided in Table I.
 Table I: Summary of commonly used abbreviations in this review.
@@
 AV Autonomous Vehicle
+ICVs Intelligent Connected Vehicles
 CAV Connected and Autonomous Vehicle
-VIC vehicle-infrastructure cooperation
 CP Cooperative or Collaborative Perception
@@
 V2I Vehicle-to-Infrastructure
-V2N Vehicle-to-Network
 I2I Infrastructure-to-Infrastructure
-RSU Roadside Unit
 ITS Intelligent Transportation System
-CAM Cooperative Awareness Message
-CPM Cooperative Perception Message
-CPS Collective Perception Service
-CAD Collaborative Autonomous Driving
-CDA Cooperative Driving Automation
-GPS Global Positioning System
-IMU Inertial Measurement Unit
 BEV Bird’s Eye View
-CNN Convolutional Neural Network
 MF Motion Forecasting
@@
 II Collaboration-Based Perception Datasets
-With the rapid development of Intelligent Connected Vehicles (ICVs) and Intelligent Transportation Systems (ITS), CP paradigms have become increasingly diverse.
-These paradigms now encompass not only pure roadside perception but also V2V, V2I, and more comprehensive V2X.
-Additionally, emerging paradigms such as I2I collaboration have recently gained attention.
+With the advancement of Intelligent Connected Vehicles (ICVs) and Intelligent Transportation Systems (ITS), CP paradigms have expanded beyond roadside perception to include V2V, V2I, V2X, and the emerging I2I frameworks.
 As illustrated in Fig.
-4, various CP architectures are now being explored.
-In this section, we classify and discuss CP datasets according to their communication paradigms, including roadside perception datasets (Section II-A), V2V datasets (Section II-B), V2I datasets (Section II-C), V2X datasets (Section II-D), and I2I datasets (Section II-E).
-Table II provides a comprehensive summary of representative datasets under each paradigm.
-These datasets play a crucial role in advancing cooperative intelligence for ICVs and ITS.
+4, these paradigms support diverse communication and fusion architectures.
+This section categorizes CP datasets by collaboration type: roadside (Section II-A), V2V (Section II-B), V2I (Section II-C), V2X (Section II-D), and I2I (Section II-E).
+Table II summarizes key datasets in each category.
+These resources are critical to driving progress in cooperative intelligence for next-generation ICVs and ITS.
 Table II: Publicly Available Datasets for Collaborative Perception
@@
 DOLPHINS [89] 2022 ACCV L, C Sim 2DOD, 3DOD 42.3k 42.3k 292.5k Link
+V2XSet-Noise [58] 2024 arXiv L, C Sim 3DOD 11.4k 11.4k - Link
 DeepAccident [50] 2024 AAAI L, C Sim 3DOD, MOT, SS, TP - 57k 285k Link
@@
 V2XPnP-Seq [92] 2024 arxiv L, C Real Perception and Prediction 208k 40k 1.45M Link
+SCOPE [93] 2024 arxiv C, L Sim 2DOD, 3DOD, SS, S2R 17k 17k 575k Link
 Mixed Signals [94] 2025 arxiv L Real 3DOD - 45.1k 240.6k Link
-SCOPE [93] 2024 arxiv C, L Sim 2DOD, 3DOD, SS, S2R 17k 17k 575k Link
+Griffin [95] 2025 arXiv C, L Sim 3DOD, MOT 275k - - Link
 I2I Rcooper [60] 2024 CVPR C, L Real 3DOD, MOT 50k 30k 30k Link
@@
 Roadside perception plays a vital role in AD and ITS by offering a global, occlusion-resilient view of traffic participants [96][97][98].
-Unlike vehicle-mounted systems, infrastructure-based sensing units—typically installed on poles, gantries, or overpasses—enable the monitoring of complex intersections, dense traffic flows, and long-range environments with enhanced spatial stability and temporal continuity.
-By supplementing on-vehicle sensing with infrastructure-level observations, roadside perception significantly improves perception robustness in scenarios involving occlusions, blind spots, or congested conditions.
+Unlike vehicle-mounted systems, infrastructure-based perception units—typically installed on poles, gantries, or overpasses—enable the monitoring of complex intersections, dense traffic flows, and long-range environments with enhanced spatial stability and temporal continuity.
+By supplementing on-vehicle perception with infrastructure-level observations, roadside perception significantly improves perception robustness in scenarios involving occlusions, blind spots, or congested conditions.
+Figure 5: Samples of roadside perception datasets. (a), (d) Ko-PER [42]; (b), (c), (e), (f) IPS300+ [75]; (g), (h) Rope3D [76]; (i) INTERACTION [44].
 A variety of representative roadside datasets have been introduced to support infrastructure-based perception research.
-Early datasets such as Ko-PER [43] and CityFlow [44] primarily target intersection environments and person or vehicle Re-ID tasks.
-INTERACTION [45] expands the behavioral diversity of roadside scenarios by incorporating multi-national data with adversarial and cooperative driving maneuvers.
-Rope3D [75] provides high-resolution 3D data along with dedicated metrics for evaluating roadside 3D object detection.
-LUMPI [76] supports multi-view analysis using synchronized multi-sensor setups, while IPS300+ [74] offers dense, long-range 3D annotations for complex urban intersections.
-TUMTraf-I [77] enhances diversity through coverage of varying lighting and weather conditions.
-For highway-oriented research, A9-Dataset [73] offers real-world, high-speed traffic scenarios, whereas RoScenes [78] features the largest perception area to date through high-density, multi-camera roadside deployment.
-H-V2X [52] extends this trend by integrating radar and camera data, supporting a range of tasks including detection, tracking, and trajectory prediction under complex highway conditions.
-The release of roadside datasets has significantly broadened the scope of perception research, enabling detailed studies on object detection, semantic segmentation, and multi-object tracking from an infrastructure-centric perspective.
-These datasets also provide a valuable foundation for future V2I CP systems by enhancing scene coverage and perceptual redundancy.
-However, several challenges remain inherent to infrastructure-based data collection.
-Such as, high deployment and maintenance costs restrict large-scale installations, particularly in highway and rural areas.
-Data distributions are often skewed toward urban environments, which limits the generalizability of trained models across diverse traffic scenarios.
-Additionally, precise temporal synchronization and spatial alignment among heterogeneous sensors remain non-trivial, especially in long-term or multi-site deployments.
+Early datasets such as Ko-PER [42] and CityFlow [43] primarily target intersections and Re-ID tasks.
+Later datasets expand coverage and functionality: INTERACTION [44] emphasizes interactive behaviors using drone-captured trajectories and semantic maps; Rope3D [76] provides top-down monocular images with 3D annotations; LUMPI [77] and IPS300+ [75] support multi-view, multi-modal fusion with rich annotations; and TUMTraf-I [78], A9-Dataset [74], RoScenes [79], and H-V2X [51] extend coverage to diverse weather, lighting, and highway scenarios.
+Fig.
+5 illustrates the diversity across Ko-PER [42], IPS300+ [75], Rope3D [76], and INTERACTION [44] in sensor modalities, viewpoints, and scene complexities.
+Ko-PER and Rope3D use monocular or stereo RGB with fixed viewpoints; Rope3D emphasizes long-range top-down detection, while Ko-PER captures lateral mid-range views.
+IPS300+ integrates LiDAR and stereo cameras for dense, occlusion-resilient 3D annotations.
+INTERACTION offers behavior-rich drone views for planning and imitation learning tasks.
+Together, they reflect trade-offs between perception complexity, spatial coverage, and task specialization.
+These datasets have expanded infrastructure-centric perception research across detection, segmentation, and tracking tasks, while also enhancing V2I systems through broader scene coverage and perceptual redundancy.
+Nonetheless, challenges persist, including deployment costs, urban-centric biases, and difficulties in sensor synchronization and alignment across deployments.
 II-B V2V Datasets
-V2V CP enhances perception range and accuracy by enabling information sharing among multiple vehicles.
-Compared to single-vehicle perception, V2V collaboration effectively mitigates blind spots caused by occlusions and compensates for the limitations of individual sensors in specific scenarios.
-This leads to improved robustness in tasks such as object detection, semantic understanding, and behavior prediction.
-Furthermore, the integration of multi-sensor data and cross-view information enables V2V systems to perform more adaptively in complex and dynamic environments.
-Therefore, the development of high-quality V2V CP datasets is of great importance for advancing research in multi-vehicle collaborative autonomous driving systems.
-In recent years, V2V datasets have continued to evolve in terms of collaborative modeling and data organization mechanisms.
-Most datasets, including OPV2V [40], MARS [82], and LUCOOP [81], adopt an offline alignment strategy that unifies data from multiple vehicles into a common coordinate system using global positioning or timestamps.
-This approach simulates idealized synchronized collaboration scenarios.
-However, it fails to account for real-world issues such as asynchronous sampling, communication latency, and localization errors during actual V2V communication.
-To address these limitations, HEAL [83] introduces asynchronous perception and temporal alignment mechanisms, while OPV2V-N [97] employs neural modeling to improve perception robustness.
-COMAP [79] adopts a task-oriented modeling approach from a multi-agent SLAM perspective.
-In terms of agent configuration, most datasets support CP among 2 to 4 vehicles.
-MARS [82] enables data collection across multiple time periods, trajectories, and viewpoints using four connected vehicles.
-OPV2V+ [80] extends viewpoint variation modeling by supporting primary view selection and perception completion.
-Regarding task coverage, existing datasets support key tasks such as object detection, tracking, Bird’s Eye View (BEV) semantic segmentation, and path planning.
-V2V-QA [84] further incorporates multi-modal large language models to support intelligent question answering and path reasoning tasks.
-IRV2V [98] and CODD [49] focus on asynchronous robustness and multi-view point cloud registration, reflecting the expanding scope of V2V data modeling in terms of collaboration granularity and semantic capabilities.
-Despite the notable progress made by V2V CP datasets in enabling multi-vehicle fusion research, several distinct challenges remain.
-For example, most existing datasets rely on offline stitching techniques for data integration and lack modeling of real communication links, transmission latency, and collaborative constraints.
-As a result, they fail to capture the uncertainty and dynamic characteristics of information exchange in real V2V scenarios, limiting deeper investigations into the essence of cooperative mechanisms.
-Looking ahead, the development of V2V datasets should focus on realistic modeling of communication processes, the design of dynamic alignment mechanisms for CP, and the construction of efficient multi-vehicle data acquisition and annotation pipelines in complex urban environments.
+V2V CP enhances perception coverage and accuracy by enabling vehicles to share information across viewpoints.
+It helps overcome occlusions and sensor limitations, supporting robust detection, semantic understanding, and motion prediction in multi-agent environments.
+High-quality datasets are thus vital for advancing V2V perception research.
+V2V datasets have evolved significantly in alignment strategies and task diversity.
+Early datasets like OPV2V [39], LUCOOP [82], and MARS [83] assume ideal synchronization via GPS or timestamps, overlooking practical challenges such as asynchrony and drift.
+Later datasets incorporate asynchronous modeling (HEAL [84]), neural alignment (OPV2V-N [99]), and SLAM-based fusion (COMAP [80]).
+OPV2V+ [81] adds view selection and perception completion.
+While most support 2–4 vehicles, MARS emphasizes trajectory diversity.
+Tasks now cover detection, tracking, Bird’s Eye View (BEV) segmentation, and planning.
+V2V-QA [85] further introduces multi-modal reasoning using LLMs; IRV2V [100] and CODD [48] focus on registration under asynchrony.
+Fig.
+6 (a–c) illustrates this progression.
+OPV2V (a) provides synthetic CARLA data for controlled and reproducible benchmarking but lacks real-world complexity.
+V2V4Real (b) contributes large-scale real LiDAR data involving dual agents in diverse scenarios.
+V2V-QA (c) extends this by introducing multi-modal reasoning tasks powered by LLMs.
+Collectively, these datasets mark a shift from basic sensor fusion toward semantically rich and cognitively informed V2V collaboration.
+Despite these advances, most V2V datasets still rely on offline stitching and neglect realistic modeling of communication delay, bandwidth constraints, and uncertainty.
+Future efforts should integrate real-time communication, dynamic alignment, and large-scale multi-agent urban scenes to support deployable cooperative strategies.
 II-C V2I Datasets
-V2I perception datasets aim to enhance AD systems by enabling collaboration between vehicles and roadside infrastructure.
-These datasets incorporate multimodal data from both vehicle-mounted and infrastructure-based sensors, offering a more comprehensive and accurate understanding of the environment.
-Representative V2I datasets exhibit distinct characteristics and collaborative strategies.
-CoopInf [48] employs eight infrastructure sensors for multi-view 3D object detection.
-DAIR-V2X-Seq [42] is the first large-scale sequential V2X dataset using data from both infrastructure and vehicles.
-HoloVIC [85] constructs holographic intersections with diverse layouts to overcome occlusion and blind spots through multisensor fusion.
-TUMTraf-V2X [87] emphasizes perception under complex lighting conditions and supports 3D detection and tracking.
-OTVIC [86] targets high-speed and noisy communication scenarios with multimodal, multi-view perception data.
+V2I perception datasets enhance AD systems by facilitating collaboration between vehicles and roadside infrastructure.
+These datasets integrate multimodal information from both vehicle-mounted and infrastructure-based sensors, enabling a more accurate and holistic understanding of the driving environment.
+Representative V2I datasets exhibit distinct characteristics and collaboration designs.
+CoopInf [47] utilizes eight infrastructure sensors to support multi-view 3D object detection.
+DAIR-V2X-Seq [41] is the first large-scale sequential V2X dataset incorporating time-aligned data from both vehicles and infrastructure.
+HoloVIC [86] creates holographic intersection scenes with diverse layouts, aiming to mitigate occlusion and blind spots through multimodal fusion.
+TUMTraf-V2X [88] focuses on perception in challenging lighting conditions and supports 3D detection and tracking tasks.
+OTVIC [87] addresses high-speed and noisy communication settings using multimodal, multi-view sensor data.
+Figure 6: Samples from representative V2V and V2I datasets. (a) OPV2V [39]; (b) V2V4Real [52]; (c) V2V-QA [85]; (d)(e)(g)(h) DAIR-V2X [35]; (f)(i) OTVIC [87].
+As illustrated in Fig.
+6 (d)–(i), DAIR-V2X ((d)(e)(g)(h)) provides synchronized vehicle-side and infrastructure-side data across various lighting and weather conditions, making it suitable for robust V2I perception under real-world deployment constraints.
+In contrast, OTVIC ((f)(i)) focuses on sparse highway scenarios with long-range detection tasks, leveraging BEV projections and trajectory-level annotations.
+The visual samples highlight the contrast in scene layout, detection distance, and data density, reflecting complementary application targets.
 While these datasets significantly contribute to improving V2I CP, challenges remain.
-Sensor synchronization is technically demanding, and many datasets offer limited scene diversity or weather coverage.
-Future efforts should prioritize building more diverse, comprehensive V2I datasets to improve perception robustness in complex and dynamic driving environments.
+Sensor synchronization is technically demanding, and many datasets offer limited scene diversity or weather coverage [101].
+Future research should prioritize more heterogeneous scene types and realistic multi-view synchronization to enhance generalization across V2I deployments.
 II-D V2X Datasets
-CP empowered by V2X technology facilitates data sharing and joint environmental understanding through communication among diverse agents, including vehicles, infrastructure, pedestrians, and other IoT devices.
-This paradigm significantly improves perception robustness and driving safety in autonomous systems.
-From the perspective of agent structure and heterogeneity, recent V2X datasets reveal an increasing trend towards modeling complex interactions among heterogeneous agents with varying sensing capabilities.
-Early works such as V2X-SIM [41] support basic cooperative tasks, yet remain limited in agent type diversity and environmental complexity.
-The DAIR-V2X series [39, 55, 56] significantly expands the agent spectrum by introducing bi-agent collaboration between vehicles and RoadSide Units (RSUs).
-These datasets illustrate asymmetric collaboration in heterogeneous settings, where infrastructure and mobile vehicles possess distinct sensor setups and motion patterns.
-To capture multi-agent temporal dynamics, V2XPnP [91] offers sequential data involving multiple cooperative vehicles, facilitating trajectory prediction in temporally correlated environments.
-DOLPHINS [88] further enhances agent diversity by simulating tri-vehicle collaboration in adverse conditions such as rain and fog, which supports evaluation of perception robustness under extreme scenarios.
-V2X-Real [54] emphasizes large-scale data collection across multiple vehicles and infrastructures, reflecting real-world V2X deployments with rich agent-level interactions.
-Adver-City [50] and DeepAccident [51] introduce environmental challenges such as snow, low visibility, and accident-prone zones, where CP across various agents becomes critical.
-Meanwhile, V2X-Radar [57] and V2X-R [58] expand the sensor modalities by incorporating 4D radar into the V2X sensing framework, highlighting the potential of cross-agent, cross-modal fusion in enhancing perception fidelity.
-Despite these advancements, V2X datasets still face challenges related to scalability, diversity, and task integration.
-Many datasets lack coverage of complex urban intersections or multi-agent scenarios with mixed mobility patterns.
-Additionally, existing benchmarks often address specific tasks in isolation, limiting their support for evaluating end-to-end CP systems across heterogeneous agents.
+V2X-based CP enables rich environmental understanding by linking vehicles, infrastructure, and other agents through communication.
+This paradigm enhances perception completeness and safety in dynamic traffic environments.
+Recent V2X datasets increasingly emphasize agent diversity and heterogeneous collaboration.
+Early datasets like V2X-SIM [40] support basic CP tasks in structured settings.
+The DAIR-V2X series [35, 54] extends to V2I collaboration with asymmetric sensor setups.
+Multi-agent dynamics are explored in V2XPnP [92] and DOLPHINS [89], which incorporate sequential reasoning and adverse weather.
+Adver-City [49], DeepAccident [50], and V2X-Real [53] improve realism and environmental diversity, while V2XSet-Noise [58] injects simulated sensor and timing noise to evaluate robustness under deployment-like conditions, and V2X-Radar [56] incorporates 4D radar for cross-modal fusion.
+Figure 7: Samples from representative V2X datasets. (a)(b)(c) V2VSet [33]; (d)(e)(f) DeepAccident [50]; (g)(h)(i) Adver-City [49].
+As shown in Fig.
+7, the three datasets reflect complementary directions in advancing V2X perception.
+V2VSet (a–c) offers clean synthetic environments for cooperative 3D detection and view fusion, but lacks dynamic risk modeling.
+DeepAccident (d–f) introduces accident-prone and motion-rich sequences to support forecasting tasks.
+Adver-City (g–i) emphasizes sensor degradation under fog, glare, and other adverse conditions.
+These samples illustrate a trade-off between controllability (V2VSet), temporal realism (DeepAccident), and visual robustness (Adver-City), jointly covering core challenges in V2X CP.
+Despite progress, V2X datasets still face scalability and integration challenges.
+Many lack coverage of complex intersection layouts or hybrid traffic patterns.
+Benchmarks often isolate perception tasks, limiting support for end-to-end CP evaluation across heterogeneous agents and adverse conditions.
 II-E I2I Datasets
-As V2X CP continues to evolve, I2I collaboration has emerged as a promising paradigm to address the limitations of single-agent roadside systems.
-While V2I, V2I and V2X datasets focus primarily on perception enhancement via VIC, their perception range is often constrained by occlusion, especially when large vehicles or urban structures obstruct the view.
-I2I CP addresses this issue by enabling multiple roadside infrastructure nodes to share and fuse data, thereby reducing blind spots and improving the robustness of perception in complex environments such as intersections and road corridors.
-Two pioneering datasets exemplify the advancement of I2I CP.
-The RCooper [60] is the first real-world, large-scale roadside CP dataset.
-It includes two typical scenarios—intersection and corridor—with infrastructure agents equipped with multi-type LiDARs and cameras.
-RCooper provides high-quality annotations, demonstrating the effectiveness of cross-infrastructure cooperation.
-In contrast, the InScope dataset [59] specifically targets occlusion-aware perception by deploying LiDARs at multiple viewpoints.
-Additionally, InScope proposes a novel quantitative metric to evaluate anti-occlusion capabilities, providing a rigorous framework for I2I perception evaluation.
-V2XPnP [91] further expands the paradigm by integrating large-scale sequential data across V2X modes, particularly I2I, offering valuable support for complex CP tasks.
-Compared to other paradigms, I2I collaboration is more effective in mitigating occlusion in dense traffic environments.
-However, Current I2I datasets face challenges in generalization, data synchronization, and scene diversity, highlighting the need for scalable deployment strategies and standardized evaluation protocols.
+As V2X CP advances, I2I collaboration emerges as a key paradigm to address the limitations of single-agent roadside systems.
+While traditional V2I and V2X enhance perception via vehicle–infrastructure communication, they remain susceptible to occlusions from static objects or large vehicles.
+I2I mitigates this by linking multiple roadside perception units via low-latency channels, enabling real-time feature or decision-level fusion.
+This improves robustness in occlusion-prone scenes such as intersections and corridors, and supports scalable, decentralized cooperative frameworks.
+Three recent datasets highlight advances in I2I perception.
+RCooper [60] is the first large-scale real-world roadside dataset, covering intersection and corridor scenes with heterogeneous infrastructure agents using diverse LiDARs and cameras.
+It provides 3D detection and tracking labels to validate cross-node fusion.
+InScope [59] emphasizes occlusion-aware perception using multi-view LiDARs and introduces a degradation metric for systematic evaluation.
+V2XPnP [92] complements these with long-sequence, multi-agent data—including I2I settings—to support temporal modeling and complex tasks.
+Figure 8: Samples of I2I collaborative perception datasets. (a)–(d) RCooper [60]; (e) InScope [59].
+Fig.
+8 compares two I2I datasets through representative samples.
+RCooper (a)–(d) features multi-angle RGB images and LiDAR point clouds from distributed infrastructure nodes, enabling long-range coverage and blind spot reduction via inter-node fusion.
+In contrast, InScope (e) employs a dual-LiDAR setup for occlusion-aware perception in dense urban scenes.
+While RCooper emphasizes spatial coverage and corridor-scale extensiveness, InScope focuses on precise detection in visually constrained environments, supported by its anti-occlusion evaluation metric.
+These datasets reflect complementary strategies in I2I perception: broad area fusion versus localized occlusion resolution.
+Compared to other paradigms, I2I collaboration proves particularly effective in mitigating occlusions within dense traffic environments.
+However, current I2I datasets still face limitations in generalization, data synchronization, and scene diversity, underscoring the need for scalable deployment strategies and standardized evaluation protocols.
+Beyond I2I, emerging collaborative paradigms such as Vehicle-to-Network (V2N), cloud-based collaboration, and Vehicle-to-Pedestrian (V2P) are drawing increasing attention.
+V2N and cloud communication facilitate global map sharing and large-scale model updates, while V2P extends perception capabilities to vulnerable road users, broadening the scope and societal impact of CP systems.
 III Simulation vs. Real-World Datasets
-Simulation and real-world datasets serve complementary roles in AD research, differing in cost, source, annotation accuracy, realism, and applicability.
-Simulation datasets, generated through virtual modeling, offer low cost, high scalability, and automatic annotations, enabling the design of diverse scenarios, including rare or hazardous conditions.
-However, the simulation-to-reality gap limits their generalization capabilities.
-Real-world datasets, collected from actual environments, provide greater complexity and realism, reflecting sensor noise and dynamic interactions, making them essential for final model validation.
-Despite their reliability, they are costly and labor-intensive to acquire and annotate, often lacking scenario diversity.
-As a result, simulation datasets are preferred for algorithm development and pre-training, while real-world datasets remain critical for evaluation and deployment [36].
+While Section II organized CP datasets according to communication paradigms such as V2V, V2I, and V2X, this section shifts perspective to the origin and fidelity of the data—specifically, the distinction between simulated and real-world datasets.
+This transition enables a deeper understanding of how dataset generation methods impact the development, evaluation, and deployment of CP systems.
+Simulation and real-world datasets play complementary roles in collaborative perception.
+Simulated datasets offer scalable, controllable environments with dense annotations, enabling rapid iteration and benchmarking.
+In contrast, real-world datasets provide authentic sensor artifacts, diverse environmental conditions, and deployment realism essential for evaluating generalization and robustness.
+However, the domain gap between synthetic and real data introduces significant challenges for transferability and real-world performance [62].
+In this section, we first review mainstream simulation platforms and representative simulated datasets, then analyze widely used real-world datasets.
+We conclude by discussing the Sim2Real gap, a key challenge in CP research, where models trained in simulation often degrade when applied to real-world environments.
+Table III summarizes the core differences between simulated and real-world datasets, highlighting their respective advantages and limitations.
 Table III: Comparison of Simulated and Real Datasets Characteristics
@@
 III-A1 Simulation Platforms
-Simulators are essential for CP research in AD, offering cost-effective, safe, and annotation-efficient alternatives to real-world data collection.
-They enable accurate generation of object attributes and automated annotations.
-Open-source platforms such as CARLA [99], SUMO [100], and AirSim [101] are widely used for data synthesis.
-CARLA supports diverse driving scenarios and multimodal sensors, including RGB, depth, radar, and LiDAR, but its LiDAR simulation has limitations in point cloud intensity and echo dropout modeling.
-SUMO provides vehicle and pedestrian trajectory simulation and can be co-simulated with CARLA to improve scene fidelity.
-The OpenCDA [102] framework integrates both platforms, offering standardized scenario management and benchmarks for perception and planning tasks.
+Simulation platforms play a pivotal role in CP research for AD, providing scalable, controllable, and annotation-rich alternatives to costly real-world data collection.
+By enabling the generation of synthetic sensor data with fine-grained ground truth, these simulators accelerate the development and validation of perception algorithms under diverse and repeatable conditions.
+Among open-source simulation platforms, CARLA [102] is one of the most widely adopted tools for AD research.
+It supports various multimodal sensors such as RGB cameras, LiDAR, radar, and depth maps, and provides configurable urban and rural scenarios.
+However, its LiDAR simulation has limitations in modeling realistic point cloud intensity and echo effects.
+SUMO [103] focuses on microscopic traffic flow and is often used to simulate vehicle and pedestrian trajectories.
+When integrated with CARLA, it enhances multi-agent interaction realism.
+To facilitate CP and V2X studies, OpenCDA [104] integrates both platforms, offering standardized scenarios, communication protocols, and benchmarks for end-to-end testing.
+AirSim [105] is another widely used simulator for vision-centric tasks, providing photorealistic environments, though it offers limited support for LiDAR and complex traffic scenes.
 III-A2 Simulation Datasets
-Simulation datasets serve as vital testbeds for CP research, offering controllable environments, automated annotations, and support for diverse or rare scenarios.
-Representative datasets are built on CARLA, SUMO, or LiDARsim platforms.
-COMAP [79] and OPV2V [40] leverage CARLA-SUMO co-simulation for multi-agent perception across intersections and urban scenes, supporting detection, segmentation, and trajectory prediction.
-Extensions such as OPV2V+ [80] and OPV2V-H [83] improve agent diversity and simulate asynchronous sensing.
-DOLPHINS [88] and Adver-City [50] incorporate adverse weather to evaluate sensor robustness.
-V2X-Sim [41], Multi-V2X [89], and DeepAccident [51] expand support for high-resolution sensors, multi-agent scheduling, and accident prediction.
-These datasets provide critical foundations for benchmarking CP algorithms and exploring Sim2Real transfer, communication modeling, and generalization across scenarios.
+Simulation datasets play a fundamental role in CP research by enabling controlled data generation, automatic annotation, and repeatable evaluation under diverse or rare conditions.
+Most simulated datasets are built upon open-source platforms such as CARLA, SUMO, and LiDARsim, which support flexible scenario modeling and multi-sensor simulation.
+Among them, COMAP [80] and OPV2V [39] employ CARLA-SUMO co-simulation to create interactive, multi-agent urban driving environments.
+These datasets support key CP tasks, including 3D object detection, semantic segmentation, and trajectory prediction.
+The OPV2V series has been extended in OPV2V+ [81] and OPV2V-H [84], introducing increased agent diversity and modeling asynchronous perception to reflect real-world temporal misalignment.
+Several recent datasets focus on robustness under challenging environmental conditions.
+DOLPHINS [89] and Adver-City [49] simulate adverse weather scenarios such as rain, fog, and snow, which are critical for evaluating sensor degradation.
+In parallel, V2X-Sim [40], Multi-V2X [90], and DeepAccident [50] introduce large-scale cooperative settings with support for high-resolution sensor streams, complex agent coordination, and accident scenario modeling.
+Together, these datasets provide robust testbeds for benchmarking CP models, studying Sim2Real transfer, and analyzing communication-aware perception strategies.
 III-B Real-World Datasets
-Real-world CP datasets provide authentic sensor data and traffic interactions essential for AD research and model validation.
+Real-world datasets are essential for validating CP systems under authentic sensor noise, traffic dynamics, and environmental complexity.
 DAIR-V2X [35] is the first large-scale multimodal real-world dataset focusing on V2I cooperation, while V2X-Seq [41] extends it with long-sequence data and traffic signal information.
@@
 HoloVIC [86] and OTVIC [87] focus on V2I setups, with OTVIC simulating real-time vehicle-RSU communication to reflect dynamic transmission challenges in infrastructure-assisted perception.
+These datasets collectively bridge the gap between simulated environments and field deployments, fostering more transferable and robust CP systems.
+III-C Sim2Real Gap
+The Sim2Real gap refers to the performance degradation observed when models trained on synthetic data are deployed in real-world environments.
+This challenge is especially pronounced in CP, where heterogeneous sensors, complex traffic dynamics, and multi-agent coordination amplify domain discrepancies.
+Although simulation platforms offer controllable and annotation-rich environments, their sensor data often lacks realistic noise, occlusions, and temporal misalignment.
+Real-world deployments, in contrast, introduce GPS drift, communication latency, and inconsistent sensor configurations, resulting in both geometric and semantic mismatches that undermine model generalization.
+To study this problem, several benchmark datasets have been developed.
+Simulated datasets such as OPV2V [39] and V2XSet [33] enable multi-agent experiments but fall short in capturing real-world uncertainties.
+Real-world datasets like V2V4Real [52] and DAIR-V2X [35] offer more realistic settings with diverse sensor setups.
+Empirical studies have quantified the Sim2Real gap across these datasets: S2R-ViT [106] reports up to 40% AP drop when transferring from OPV2V to V2V4Real under deployment noise; CoPEFT [107] improves AP@0.7 from 21.7% to 41.8% on DAIR-V2X via parameter-efficient tuning.
+DUSA [108] and CUDA-X [109] achieve 5–20% AP gains through domain-alignment modules.
+V2X-ReaLO [110] further reveals a 13.2 mAP drop during online deployment, highlighting the impact of latency and pose noise.
+Collectively, these findings show that the Sim2Real gap in CP stems from both data distribution shifts and system-level constraints.
+Addressing it requires more realistic simulation environments, domain adaptation techniques, and hybrid training protocols that bridge synthetic and real domains.
+Future research should also prioritize time-aware, uncertainty-aware modeling and develop standardized benchmarks for cross-domain evaluation to ensure robust real-world deployment.
 IV Cooperative perception Scenario Datasets
-CP datasets can be categorized according to different traffic environments to address the perception needs of AD systems in various scenarios.
-In this section, we classify them into three types: intersection, urban street, and highway scenarios.
+Building on the discussion of data sources in Section III, we now focus on the deployment contexts of CP datasets.
+Scenario-specific datasets are essential for evaluating system performance under diverse traffic conditions and supporting targeted algorithm development.
+This section categorizes CP datasets into three representative scenarios: intersections, urban streets, and highways.
+These settings differ in traffic complexity, agent density, and occlusion patterns, shaping dataset design and system requirements.
 Fig.
-5 shows the distribution of representative datasets across these categories.
-This chapter discusses the key datasets in each scenario and their characteristics, providing essential data support for CP research.
+9 illustrates the distribution of representative datasets.
+For each category, we review key datasets, emphasizing their design motivations, sensor configurations, and task relevance.
 Figure 9: A Venn diagram showing the distribution of CP datasets across intersection, urban street, and highway scenarios. The overlapping areas indicate datasets applicable to multiple scenarios.
@@
 The following datasets are among the most representative in this category.
-Ko-PER [43] is one of the earliest datasets employing fixed LiDAR and cameras to monitor intersection dynamics.
-IPS300+ [74] improves spatial resolution with high-density 3D annotations, while LUMPI [76] introduces multi-view data under varying weather and lighting.
-TUMTraf-I [77] further targets annotation precision in adverse conditions, such as night and rain.
-HoloVIC [85] designs diverse holographic intersections to explore sensor placement strategies for CP.
-Several datasets incorporate intersection scenarios within broader V2X contexts.
-DAIR-V2X [39] supports V2I perception in urban crossings; DOLPHINS [88] includes intersection and T-junctions with weather variations; DeepAccident [51] focuses on accident prediction at intersections.
-V2XPnP [91] offers 163 intersection scenarios, supporting trajectory prediction and object tracking.
-These intersection datasets collectively emphasize high-precision perception in dense, occlusion-prone environments through multi-view sensing and cross-agent fusion.
-However, challenges such as dynamic traffic variability, sensor blind zones remain.
-These intersection datasets collectively emphasize high-precision perception in dense, occlusion-prone environments through multi-view sensing and cross-agent fusion.
+Ko-PER [42] is one of the earliest datasets using fixed LiDAR and cameras to observe intersection dynamics.
+IPS300+ [75] enhances resolution with dense 3D annotations, while LUMPI [77] introduces multi-view data under diverse weather and lighting.
+TUMTraf-I [78] focuses on annotation precision in night and rain conditions.
+HoloVIC [86] designs holographic intersections to study sensor deployment.
+Several V2X datasets also incorporate intersection scenarios: DAIR-V2X [35] supports V2I in urban crossings; DOLPHINS [89] adds weather-varied intersections; DeepAccident [50] centers on accident prediction; V2XPnP [92] includes 163 intersections for trajectory prediction and tracking.
+These datasets emphasize high-precision perception in occlusion-prone intersections via multi-view perception and cross-agent fusion, though challenges like traffic variability and sensor blind spots remain.
 IV-B Urban Street Datasets
-Urban street CP datasets span V2V, V2I, and V2X paradigms, supporting tasks such as object detection, tracking, and trajectory prediction.
-CityFlow [44] offers multi-camera tracking data across arterials and residential areas.
-LUCOOP [81] collects time-synchronized multimodal data from three Connected Autonomous Vehicles (CAVs), including 3D point clouds, city models, and UWB ranging data.
-MARS [82] further extends to four CAVs under diverse weather, enhancing perception stability and agent collaboration.
-Several datasets strengthen roadside sensing capabilities.
-RCooper [60] provides the first large-scale real-world roadside dataset for multi-sensor infrastructure perception.
-RoScenes [78] delivers dense multi-view 3D data in complex urban traffic, supporting BEV perception.
-V2XPnP [91] captures diverse urban road scenes with multi-agent trajectories and detailed interactions.
-Together, these datasets address complex urban traffic environments, offering comprehensive data support for perception, fusion, and prediction in multi-agent V2X systems.
+Urban street CP datasets span V2V, V2I, and V2X paradigms, enabling tasks such as object detection, tracking, and trajectory prediction.
+CityFlow [43] provides multi-camera tracking across arterial and residential roads.
+LUCOOP [82] uses synchronized multimodal data from three Connected and Autonomous Vehicles (CAVs), including 3D point clouds, city models, and UWB ranging.
+MARS [83] extends to four CAVs under diverse weather, enhancing collaborative robustness.
+Roadside-focused datasets also emerge: RCooper [60] is the first large-scale real-world roadside CP dataset; RoScenes [79] offers dense multi-view 3D urban scenes for BEV perception; V2XPnP [92] captures multi-agent urban trajectories and interactions.
+Collectively, these datasets provide comprehensive urban traffic data for multi-agent perception, fusion, and prediction.
 IV-C Highway Datasets
 With the growing deployment of ICVs on highways, CP datasets for highway scenarios have become increasingly important for traffic management and AD safety.
-Highway environments demand long-range detection, lane-level perception, and robust environmental sensing, especially under high-speed conditions where precise synchronization is critical.
+Highway environments demand long-range detection, lane-level perception, and robust environmental perception, especially under high-speed conditions where precise synchronization is critical.
 Although limited in number, highway CP datasets are gradually expanding.
@@
 V Datasets by Sensor Modality
-Sensor modality diversity plays a pivotal role in enhancing the robustness of CP systems, particularly in complex and dynamic environments [103][104].
-As illustrated in Fig.
-6, typical sensors employed in CP include cameras, LiDAR, and Radar, each corresponding to distinct perception modalities such as RGB images, dense 3D point clouds, and sparse Radar point clouds.
-The respective strengths and weaknesses of these sensors are summarized in Table IV, providing a comprehensive overview of their capabilities under varying operational conditions.
-This section categorizes existing CP datasets based on their acquisition modalities, including image-only, LiDAR-only, and multimodal datasets.
+Sensor modality diversity is critical for enhancing the robustness of CP systems in complex, dynamic environments [111][112].
+As shown in Fig.
+10, CP typically employs cameras, LiDAR, and Radar, corresponding to RGB images, dense 3D point clouds, and sparse Radar reflections.
+Image-only datasets are cost-effective and semantically rich but suffer from limited depth perception and sensitivity to lighting.
+LiDAR-only datasets provide accurate 3D localization and perform well in low-light conditions but lack semantic information and are costly to deploy.
+Multimodal datasets, particularly those integrating cameras and LiDAR, are prevalent due to their complementary strengths.
+Radar-integrated datasets have gained traction for their robustness in adverse weather [56].
+Table IV summarizes the trade-offs among these modalities in terms of resolution, cost, weather resilience, and perception reliability.
+This section categorizes CP datasets based on sensor modality into image-only, LiDAR-only, and multimodal types.
 Figure 10: Representative sensor types and corresponding perception modalities.
@@
 All data samples are derived from the V2X-Radar [56] dataset.
-Image-only datasets, while cost-effective and semantically rich, are limited by poor depth estimation.
-LiDAR-only datasets offer precise 3D localization but are constrained by high sensor cost and limited semantic information.
-Multimodal datasets, especially those combining camera and LiDAR, dominate recent research due to their complementary strengths.
-Moreover, Radar-integrated datasets have gained attention for their resilience to adverse weather, marking a growing trend in sensor fusion for reliable perception [57].
 Table IV: The Main Advantages and Disadvantages of Cameras, LiDAR, and Radar
@@
 • Typically expensive.
-• Performance degrades in rain, snow, and is affected by smoke, dust, and other weather conditions.
+• Performance degrades in rain, snow, and affected by smoke, dust, and weather conditions.
 • Large data volume, complex processing.
@@
 WIBAM [113], captured from infrastructure-mounted cameras, focuses on 2D/3D object detection.
-RoScenes [78], the largest multi-view roadside dataset to date, targets vision-based BEV perception, addressing challenges in multi-camera fusion and dense traffic scenes.
+RoScenes [79], the largest multi-view roadside dataset to date, objects vision-based BEV perception, addressing challenges in multi-camera fusion and dense traffic scenes.
 Despite their effectiveness in visual tasks, image-only datasets have limited utility in high-precision localization and CP.
@@
 Multi-modal datasets, particularly those fusing LiDAR and cameras, are central to CP research.
-By combining LiDAR’s precise 3D spatial sensing with camera-derived semantic richness, these datasets improve environmental understanding and robustness [106][107].
+By combining LiDAR’s precise 3D spatial perception with camera-derived semantic richness, these datasets improve environmental understanding and robustness [114][115].
 Recently, the integration of radar has further enhanced perception under adverse weather and for long-range detection.
 This section summarizes key LiDAR-Camera and LiDAR-Radar-Camera datasets.
-LiDAR-Camera datasets include both infrastructure-based and cooperative datasets.
-Fixed-infrastructure datasets such as Ko-PER [43], IPS300+ [74], Rope3D [75], A9-Dataset [73], and LUMPI [76] use roadside LiDAR and cameras to produce accurate 3D annotations.
-H-V2X [52], RCooper [60], and TUMTraf-I [77] extend perception to challenging conditions such as rain and night.
-UAV and traffic camera datasets like INTERACTION [45] and BAAI-VANJEE [108] support interaction modeling and driving behavior analysis.
-CP datasets such as OPV2V [40], IRV2V [98], V2V4Real [53], LUCOOP [81], and MARS [82] collect multi-agent LiDAR-Camera data in both real and simulated settings.
-OPV2V-N [97] enhances camera robustness, while V2V-QA [84] introduces LLM-based collaborative tasks.
-CoPerception-UAV [35] supplements V2V data with aerial views.
-In V2X contexts, DAIR-V2X-C [39], V2X-Seq [42], HoloVIC [85], and OTVIC [86] support multi-modal detection, tracking, and planning.
-Simulated datasets such as V2X-Sim 2.0 [41], V2XSet [34], DOLPHINS [88], DeepAccident [51], V2X-Real [54], and Multi-V2X [89] further address diverse tasks from accident prediction to V2X data augmentation.
-LiDAR-Radar-Camera datasets such as T & J [46], OPV2V+ [80], OPV2V-H [83], and V2X-Radar [57] enhance CP under adverse conditions by leveraging radar’s robustness to weather and long-range sensing.
-In summary, LiDAR-Camera datasets remain the cornerstone of CP research, while LiDAR-Radar-Camera fusion introduces new capabilities for perception in challenging environments.
+LiDAR-Camera datasets are foundational to CP research, covering both infrastructure-based and collaborative scenarios.
+Ko-PER [42], IPS300+ [75], Rope3D [76], and A9-Dataset [74] use roadside sensors for accurate 3D annotation, while H-V2X [51], RCooper [60], and TUMTraf-I [78] address challenging conditions such as rain and nighttime.
+INTERACTION [44] supports behavioral analysis from UAV perspectives.
+Collaborative datasets like OPV2V [39], IRV2V [100], and V2V4Real [52] enable multi-agent perception.
+OPV2V-N [99] improves image robustness, and V2V-QA [85] introduces LLM-based reasoning.
+In V2X contexts, DAIR-V2X-C [35], V2X-Seq [41], and HoloVIC [86] support detection and planning.
+Simulated datasets such as V2X-Sim 2.0 [40], V2XSet [33], DOLPHINS [89], DeepAccident [50], V2X-Real [53], and Multi-V2X [90] address a wide range of cooperative tasks.
+LiDAR-Radar-Camera datasets, including OPV2V+ [81], OPV2V-H [84], and V2X-Radar [56], enhance robustness under adverse conditions.
+Despite progress, challenges remain in aligning heterogeneous data and ensuring scalability across modalities—future work should emphasize cross-domain generalization and efficient multimodal fusion strategies.
 VI Perception Tasks and Metrics
@@
 This chapter categorizes them by task type, including object detection, semantic segmentation, object tracking, and trajectory prediction.
-Object detection focuses on 3D localization of targets, while semantic segmentation enables pixel-level scene understanding.
+Object detection focuses on 3D localization of objects, while semantic segmentation enables pixel-level scene understanding.
 Object tracking and trajectory prediction involve continuous monitoring and future motion inference of dynamic agents.
@@
 VI-A1 Datasets for Object Detection
-Multiple datasets support V2X object detection across different collaboration modes and sensing modalities.
-DAIR-V2X-C [39] provides a real-world V2I benchmark for multi-modal detection.
-V2X-Sim 2.0 [41], offers 3D detection data for cooperative agents.
-HoloVIC [85] constructs holographic intersections for multi-view 3D sensing.
-V2X-Real [54] supports both V2I and V2V detection at scale, while V2V4Real [53] focuses on diverse road types for collaborative 3D detection.
-OPV2V [40] and OPV2V+ [80] simulate multi-agent cooperation, with the latter offering improved detection benchmarks.
-For roadside perception, RCooper [60] enhances infrastructure-based 3D detection under occlusions using multiple sensor layouts.
-Rope3D [75], while not a V2X dataset, provides monocular 3D detection benchmarks that benefit cooperative vision research.
-Overall, these datasets span real and simulated settings, support V2I and V2V collaboration, and offer diverse modalities, forming a solid foundation for V2X object detection studies.
+Numerous datasets have been developed to support V2X object detection across varying collaboration paradigms and perception modalities.
+Real-world datasets such as DAIR-V2X-C [35], V2X-Real [53], and V2V4Real [52] provide large-scale, multi-modal annotations for V2I and V2V scenarios.
+Simulated datasets like OPV2V [39], OPV2V+ [81], and V2X-Sim 2.0 [40] offer controlled environments for benchmarking.
+HoloVIC [86] and RCooper [60] emphasize occlusion-aware perception via infrastructure-side perception.
+Although Rope3D [76] is not V2X-specific, it contributes monocular 3D benchmarks relevant to cooperative perception.
+Collectively, these datasets form a diverse foundation for V2X detection research.
+As shown in Fig.
+11, recent methods demonstrate improved performance across benchmarks.
+On V2V4Real, ERMVP [118] produces more complete and better-aligned bounding boxes than V2VNet [46] and CoBEVT [119], especially in dense scenes.
+On DAIR-V2X, V2X-ViTv2 [120] detects more dynamic vehicles with lower spatial error than V2VNet and Where2comm [34], aided by transformer-based fusion.
+On OPV2V, MKD-Cooper [121] surpasses DiscoNet [122] and V2X-ViT [33] in producing orientation-aware results with fewer false positives on complex road geometries.
+These comparisons highlight the impact of advanced fusion strategies in both real and simulated domains.
+Figure 11: Qualitative detection results across three datasets: (a–c) V2V4Real with V2VNet, CoBEVT, and ERMVP [118]; (d–f) DAIR-V2X with V2VNet, Where2comm, and V2X-ViTv2 [120]; (g–i) OPV2V with OPV2V, DiscoNet, and MKD-Cooper [121].
+Green boxes show ground truth; red/orange boxes indicate predictions.
+ERMVP and MKD-Cooper yield more accurate and robust results in complex scenes.
+Images reproduced from the original publications with proper citation.
 VI-A2 Metrics for Object Detection
@@
 Commonly used metrics include AP3D (based on 3D IoU) and APBEV (based on BEV IoU), as introduced by the KITTI benchmark [36].
-Most CP datasets adopt AP or mAP as core metrics.
-RCooper [60] reports AP at multiple IoU thresholds.
-OPV2V [40] and OPV2V+ [80] evaluate vehicle detection at IoU@0.5 and IoU@0.7, with performance also measured within communication ranges.
-V2X-Real [54] employs both AP and mAP, using lower IoU thresholds (0.3/0.5) to account for varying object sizes.
-V2X-Sim 2.0 [41] focuses on BEV-based AP at multiple thresholds.
-Some datasets integrate communication cost into evaluation.
-V2V4Real [53] reports AP@0.5 and AP@0.7, and introduces Average MegaByte (AM) to quantify transmission overhead under synchronous and asynchronous conditions.
-DAIR-V2X [39] incorporates communication efficiency and limits evaluation to objects within the ego vehicle’s ROI.
-Rope3D [75] enhances AP-based evaluation with additional metrics such as Average Ground Center Similarity (ACS), Average Orientation Similarity (AOS), Average Area Similarity (AAS), and Average Four Ground Points Distance (AGD) to assess localization, orientation, and shape accuracy.
-In addition to metric design, a range of benchmark methods have been evaluated across typical CP datasets.
-Table V summarizes the performance of representative algorithms on DAIR-V2X, V2V4Real, OPV2V, and V2XSet, offering insights into algorithm advancements and dataset challenges.
-In summary, AP and IoU remain dominant evaluation standards in CP datasets, while extended metrics—such as BEV AP, communication cost, and asynchronous performance—offer deeper insights into real-world constraints.
-However, AP treats all missed detections equally, regardless of safety relevance.
-Future research should explore safety-aware evaluation frameworks tailored to CP tasks.
+Most datasets use AP or mAP as core metrics.
+RCooper [60] reports AP under multiple IoUs; OPV2V [39] and OPV2V+ [81] assess detection at 0.5 and 0.7 IoU, including within communication range.
+V2X-Real [53] adopts lower thresholds (0.3/0.5) for scale variance.
+V2X-Sim 2.0 [40] emphasizes BEV-based AP.
+V2V4Real [52] introduces Average MegaByte (AM) to quantify communication cost under sync/async settings, while DAIR-V2X [35] evaluates within ego-ROI under bandwidth constraints.
+Rope3D [76] supplements AP with metrics like ACS, AOS, AAS, and AGD for comprehensive localization and shape assessment.
+Beyond metrics, representative methods are benchmarked on datasets such as DAIR-V2X, V2V4Real, OPV2V, and V2XSet, as summarized in Table V, revealing advances and ongoing challenges.
+In conclusion, AP and IoU remain central for evaluating CP detection, while extended metrics—e.g., BEV AP, communication overhead, and asynchronous accuracy—offer deeper realism.
+Future work should incorporate safety-aware evaluation to prioritize critical perception failures.
 Table V: Performance Comparison of Representative Methods on Benchmark CP Datasets. All values are reported as AP@0.5 / AP@0.7. “-” indicates results not reported in the corresponding paper.
@@
 VI-B1 Datasets for Semantic Segmentation
-Several datasets support semantic segmentation in CP tasks.
-V2X-Sim 2.0 [41] provides multi-agent LiDAR segmentation with dynamic obstacle, lane, and BEV semantic annotations.
-OPV2V [40] supports BEV semantic segmentation using LiDAR-camera fusion, while OPV2V+ [80] enhances cross-agent temporal alignment.
-OPV2V-N [97] focuses on improving image segmentation robustness.
-COMAP [79] offers BEV semantic segmentation in simulated urban scenarios.
-DeepAccident [51] provides accident scene segmentation with fine-grained semantic labels.
-DOLPHINS [88] supports V2X BEV segmentation under varied weather.
-Adver-City [50] collects segmentation data under extreme weather using dedicated semantic cameras.
-Despite these advances, semantic segmentation in CP still faces several challenges, including inconsistent semantic representations across agents, and reduced robustness in adverse weather or sparse data conditions.
-Moreover, current datasets are limited in diversity and scale, especially for real-world multi-agent segmentation benchmarks.
-Future trends will emphasize the construction of large-scale real-world V2X segmentation datasets, cross-agent temporal consistency, and robustness under extreme conditions to support reliable and generalizable perception in complex collaborative scenarios.
+Several datasets support semantic segmentation for CP.
+V2X-Sim 2.0 [40] provides multi-agent LiDAR segmentation with BEV and dynamic obstacle annotations.
+OPV2V [39] enables BEV segmentation via LiDAR-camera fusion, while OPV2V+ [81] and OPV2V-N [99] improve temporal alignment and image-based robustness, respectively.
+COMAP [80] offers simulated urban BEV annotations.
+DeepAccident [50] provides fine-grained accident scene labels.
+DOLPHINS [89] and Adver-City [49] support segmentation under adverse conditions using diverse sensor setups.
+Challenges remain, including inconsistent semantic labeling across agents and limited robustness under sparse or extreme conditions.
+Real-world datasets are also limited in scale and diversity.
+Future efforts should focus on large-scale, real-world, multi-agent segmentation benchmarks with strong cross-agent consistency and resilience to weather and domain shifts.
 VI-B2 Metrics for Semantic Segmentation
-Semantic segmentation in CP tasks is commonly evaluated using Pixel Accuracy (PA), mean Pixel Accuracy (mPA), Intersection over Union (IoU), and mean IoU (mIoU) [120, 121].
-PA measures the ratio of correctly classified pixels, mPA averages PA across classes to address imbalance, and mIoU reflects overall segmentation accuracy by averaging per-class IoU.
-In CP scenarios, evaluation extends beyond single-agent performance to assess cross-agent consistency and modality alignment.
-V2X-Sim 2.0 [41] evaluates BEV semantic segmentation across categories such as lanes, vehicles, and buildings using mIoU.
-OPV2V-N [97] focuses on map-level segmentation (drivable area, lanes, vehicles), using IoU as the core metric.
-Effective CP segmentation evaluation requires traditional metrics and additional measures such as temporal stability, cross-modal IoU, and multi-agent coherence.
+Semantic segmentation in CP is typically evaluated using Pixel Accuracy (PA), mean PA (mPA), Intersection over Union (IoU), and mean IoU (mIoU) [130][131].
+PA measures the proportion of correctly classified pixels, while mPA and mIoU address class imbalance and overall accuracy.
+In CP settings, evaluation extends to inter-agent consistency and modality fusion.
+V2X-Sim 2.0 [40] reports mIoU for semantic categories such as lanes and vehicles.
+OPV2V-N [99] focuses on map-level segmentation using IoU.
+Advanced evaluation should also consider temporal stability, cross-modal consistency, and multi-agent alignment.
 VI-C Tracking and Prediction
-VI-C1 Datasets for Tracking and Prediction
 Object tracking and trajectory prediction focus on modeling the temporal dynamics of moving agents and are essential for AD.
 CP improves tracking accuracy and trajectory prediction by enabling inter-agent information sharing, particularly under occlusion and complex traffic conditions.
-Representative datasets include V2X-Seq [42], V2XPnP-Seq [91], and DAIR-V2X-Traj [56].
-V2X-Seq offers long-sequence data for both V2V and V2I scenarios, while V2XPnP-Seq enhances temporal consistency and multi-view target association.
-DAIR-V2X-Traj focuses on V2I trajectory prediction with high-precision annotations.
-DeepAccident [51] addresses accident prediction by providing trajectory data in critical traffic scenarios.In urban-scale environments, CityFlow [44] supports cross-camera object tracking for traffic monitoring, and INTERACTION [45] provides detailed vehicle trajectories for interactive behavior modeling.
-RCooper [60] employs roadside sensors to support tracking in dense traffic.
-OPV2V [40] and V2V4Real [53] benchmark V2V tracking and trajectory prediction in simulated and real-world settings, respectively.
-V2X-Sim 2.0 [41] supports multi-agent tracking and prediction using BEV-based annotations.
-Overall, CP datasets for tracking and prediction span V2V, V2I, and roadside sensing paradigms, providing diverse benchmarks for multi-agent temporal perception and forecasting in AD systems.
+VI-C1 Datasets for Tracking and Prediction
+Representative datasets for tracking and prediction span various CP paradigms.
+V2X-Seq [41] and V2XPnP-Seq [92] support long-sequence V2V/V2I perception with improved temporal consistency and multi-view association.
+DAIR-V2X-Traj [55] provides accurate V2I trajectory annotations, while DeepAccident [50] focuses on forecasting in safety-critical scenarios.
+Urban-scale datasets like CityFlow [43] and INTERACTION [44] enable cross-camera tracking and behavior modeling.
+RCooper [60] captures dense roadside tracking scenes, and OPV2V [39] and V2V4Real [52] benchmark V2V tracking in simulation and real-world contexts.
+V2X-Sim 2.0 [40] supports BEV-based multi-agent prediction.
+Collectively, these datasets provide diverse benchmarks for evaluating temporal perception across V2V, V2I, and infrastructure-based CP.
 VI-C2 Metrics for Tracking and Prediction
-Object tracking and trajectory prediction are central to CP, requiring precise evaluation of detection accuracy, tracking stability, and motion forecasting.
-Standard metrics have been developed to meet these requirements, with extensions introduced to address the complexities of multi-agent V2X scenarios.
-For tracking tasks, the CLEAR MOT framework is widely adopted, incorporating Multiple Object Tracking Accuracy (MOTA), Precision (MOTP), and Identity F1 Score (IDF1).
-MOTA considers False Positives (FP), False Negatives (FN), and Identity Switches (IDSW), while MOTP measures average IoU between predicted and ground-truth trajectories.
-IDF1 evaluates consistency in identity association.
-Several datasets extend this framework.
-RCooper [60] and V2V4Real [53] incorporate Average MOTA (AMOTA), Average MOTP (AMOTP), and scaled AMOTA (sAMOTA) to assess tracking under varied object sizes and detection thresholds.
-DAIR-V2X-Traj [56] and V2X-Seq [42] also report MOTA, MOTP, and ID Switches for comprehensive analysis.
-To better reflect multi-agent collaboration, V2X-Sim 2.0 [41] introduces Higher Order Tracking Accuracy (HOTA), which jointly evaluates detection, association, and temporal stability.
-It also uses Byte Per Second (BPS) to quantify transmission cost, enabling analysis of the trade-off between bandwidth and accuracy.
-Trajectory prediction tasks primarily use Average Displacement Error (ADE) and Final Displacement Error (FDE).
-ADE measures the mean offset over the entire predicted trajectory, while FDE focuses on the final point prediction accuracy.
-Datasets like INTERACTION [45] and V2X-Seq [42] adopt these metrics, with V2X-Seq further introducing Missing Rate (MR) to capture trajectory coverage completeness.
-In accident prediction, DeepAccident [51] proposes Volumetric Prediction Quality (VPQ) to jointly evaluate object detection and trajectory forecasting.
-It also introduces Accident Prediction Accuracy (APA), based on True/False Positives and Negatives, offering a safety-aware evaluation scheme for high-risk scenarios.
-Overall, MOTA, MOTP, ADE, and FDE remain core metrics for tracking and prediction, while newer metrics like HOTA, BPS, VPQ, and APA enable more comprehensive evaluations tailored to V2X CP.
-These frameworks offer critical tools for benchmarking CP performance and guiding future system optimization.
+Tracking and prediction in CP rely on metrics that evaluate detection accuracy, identity consistency, and MF.
+Standard CLEAR MOT metrics—MOTA, MOTP, and IDF1—remain widely used.
+Datasets like RCooper [60] and V2V4Real [52] extend these with AMOTA and sAMOTA for scale-aware evaluation.
+DAIR-V2X-Traj [55] and V2X-Seq [41] report similar metrics, while V2X-Sim 2.0 [40] introduces HOTA and BPS to account for association quality and communication cost.
+Trajectory prediction is typically assessed using ADE and FDE, as adopted by INTERACTION [44] and V2X-Seq [41], with the latter also reporting Missing Rate (MR).
+DeepAccident [50] introduces VPQ and APA to evaluate detection-prediction synergy and accident classification.
+While MOTA and ADE remain core metrics, scenario-aware metrics like HOTA and VPQ offer deeper performance insights.
 VII Comprehensive Analysis of Collaborative Perception Datasets
@@
 VII-A Availability and Access Methods of Datasets
-Dataset accessibility plays a critical role in promoting research progress in CP [123] .
-Publicly available datasets reduce annotation costs and enable reproducibility, fostering innovation and collaboration.
-To assess the openness of current datasets, we categorize existing CP datasets based on their access methods.
-First, a large portion of datasets can be directly downloaded from official websites, offering high accessibility without requiring registration or application.
-These include Ko-PER [43], CityFlow [44], LUMPI [76], TUMTraf-I [77], RoScenes [78], OPV2V [40], OPV2V+ [80], V2V4Real [53], LUCOOP [81], TUMTraf V2X [87], V2X-Sim 2.0 [41], DeepAccident [51], V2X-Real [54], Multi-V2X [89], Adver-City [50], MARS [82], Mixed Signals [92], and SCOPE [93].
-In addition, some datasets are hosted on GitHub, often accompanied by source code, benchmark tools, or baseline implementations.
-Representative examples include V2V-Sim [47], OPV2V-H [83], CoPerception-UAV [35], CoopInf [48], V2XSet [34], and V2X-R [58], which facilitate rapid experimentation and method replication.
-Meanwhile, several datasets are distributed via cloud platforms such as Baidu Netdisk, Google Drive, or Alibaba Cloud.
-Access typically requires a request or registration process.
-These include H-V2X [52], DAIR-V2X-C [39], V2X-Seq [42], OTVIC [86], V2XReid [55], V2X-Radar [57], WHALES [90], and V2XPnP-Seq [91].
-Furthermore, a number of datasets require formal application procedures, often involving submission of research proposals or institutional endorsements.
-Examples include BAAI-VANJEE [108], Rope3D [75], A9-Dataset [73], HoloVIC [85], V2V-QA [84], CODD [49], DAIR-V2X-Traj [56], INTERACTION [45], IPS300+ [74], DOLPHINS [88], InScope [59], and RCooper [60].
-Finally, a few datasets do not publicly disclose their access procedures, possibly due to ongoing internal usage, limited availability, or institutional restrictions.
-These include WIBAM [105], T & J [46], IRV2V [98], OPV2V-N [97], CARTI [124], and V2X-DSI [125].
-In summary, while many CP datasets are openly accessible via official websites, GitHub, or cloud storage, a significant portion remains gated behind application-based processes or lacks clear documentation.
-Improving data openness, unifying access procedures, and providing transparent distribution mechanisms will be essential to advancing CP research.
+Dataset accessibility plays a crucial role in advancing CP research [133].
+Public availability reduces annotation costs, enhances reproducibility, and promotes innovation across the community.
+To evaluate current openness, we categorize CP datasets by their access methods.
+Many datasets, including Ko-PER [42], CityFlow [43], OPV2V [39], V2V4Real [52], and DeepAccident [50], support direct download from official websites without registration.
+Others, such as V2V-Sim [46], V2XSet [33], and CoopInf [47], are hosted on GitHub alongside code and benchmarks, facilitating reproducible research.
+Datasets like DAIR-V2X-C [35], OTVIC [87], and WHALES [91] require requests through cloud storage services.
+Some datasets, including Rope3D [76], INTERACTION [44], and RCooper [60], involve formal application processes and institutional approval.
+A few others, such as IRV2V [100] and OPV2V-N [99], lack clear access instructions.
+Detailed access types and download links are summarized in Table II.
+In summary, while many CP datasets are readily available, a notable portion remains behind restricted access or insufficient documentation.
+Enhancing openness, standardizing access procedures, and improving transparency in dataset distribution will be key to accelerating progress in CP research.
 VII-B Comprehensive Quality Analysis of Datasets
@@
 As summarized in Table II and illustrated qualitatively in Fig.
-7, we conduct a multi-dimensional analysis across six key aspects: data frame volume, annotation richness, agent diversity, scenario complexity, sensor modalities, and supported tasks.
+12, we conduct a multi-dimensional analysis across six key aspects: data frame volume, annotation richness, agent diversity, scenario complexity, sensor modalities, and supported tasks.
 These dimensions reflect both dataset scalability and task coverage, providing a unified view for comparative assessment.
@@
 12 further visualize dataset strengths and limitations across these dimensions within categories such as V2V, V2I, and roadside perception.
-Overall, high-quality datasets exhibit complementary characteristics: real-world datasets like DAIR-V2X [39] and V2X-Real [54] emphasize fidelity and generalization, while simulation-based datasets like OPV2V+ [80] and Multi-V2X [89] excel in scale, flexibility, and controlled benchmarking.
-The increasing inclusion of diverse modalities, rich annotations, and comprehensive task coverage highlights a clear trend toward more holistic, multi-agent CP datasets.
-Future research would benefit from standardized benchmarking protocols across these high-quality datasets to promote reproducibility and cross-domain evaluation.
+High-quality CP datasets exhibit complementary strengths: real-world datasets such as DAIR-V2X [35] and V2X-Real [53] prioritize realism and generalization, while simulation datasets like OPV2V+ [81] and Multi-V2X [90] offer scalability and control.
+The trend toward multimodal data, rich annotations, and task diversity suggests a shift to more holistic, multi-agent CP benchmarks.
+Standardized evaluation protocols are needed to enhance reproducibility and cross-domain comparability.
 VII-C Summary of High-Influence Datasets
 In the field of CP, several milestone datasets have gained prominence. This section summarizes key high-impact datasets across infrastructure-based, V2I, V2V, and V2X CP domains.
-High-impact datasets are typically defined by high citation rates and widespread use in benchmark evaluations.
-For infrastructure-based perception, CityFlow [44], INTERACTION [45], and Rope3D [75] are widely used in multi-camera tracking and urban scene understanding.
-Representative methods such as BEVHeight [126], BEVHeight++ [127], and ViT-CoMer [128] have been developed using these datasets.
-In V2V perception, OPV2V [40] is one of the most cited simulation datasets, supporting algorithms like Where2comm [35], CoBEVT [114], and OPV2V-H [83].
-In the V2I domain, DAIR-V2X [39] and its extensions (e.g., DAIR-V2X-Seq [42], DAIR-V2X-Traj [56]) provide large-scale real-world data with high-quality multimodal annotations, supporting tasks such as detection, tracking, and motion forecasting.
-These datasets underpin studies including CoCa3D [80], HM-ViT [129], BEV-V2X [130], and How2comm [131].
-In the V2X category, V2X-Sim [41] supports BEV detection and semantic segmentation, and is frequently adopted by methods such as VoxFormer [132] and V2X-ViT [34].
+High-impact datasets are characterized by high citation rates and widespread benchmark adoption.
+In infrastructure-based perception, CityFlow [43], INTERACTION [44], and Rope3D [76] are widely used for multi-camera tracking and urban understanding, supporting methods like BEVHeight [134], BEVHeight++ [135], and ViT-CoMer [136].
+In V2V, OPV2V [39] enables numerous algorithms including Where2comm [34], CoBEVT [119], and OPV2V-H [84].
+DAIR-V2X [35] and its variants [41, 55] support V2I tasks like detection and tracking, used in CoCa3D [81], HM-ViT [137], and BEV-V2X [138].
+For V2X, V2X-Sim [40] facilitates BEV detection and segmentation, widely adopted by VoxFormer [139] and V2X-ViT [33].
 In summary, datasets like CityFlow, INTERACTION, Rope3D, OPV2V, DAIR-V2X, and V2X-Sim have become cornerstones of CP research due to their broad adoption and benchmark availability.
@@
 Future research should explore generating extreme scenario data through simulation, combined with real-world data augmentation, to enhance model responsiveness in hazardous situations.
-Moreover, anomaly detection and Out-Of-Distribution (OOD) analysis can help identify extreme scenarios, improving the safety and stability of CP systems.
+Moreover, anomaly detection and Out-Of-Distribution analysis can help identify extreme scenarios, improving the safety and stability of CP systems.
 VIII-A5 Geographic Diversity in Datasets
@@
 VIII-D Domain Adaptation and Generalization
-Domain adaptation remains a key challenge in improving dataset generalization.
-Due to differences in data collection environments, sensor configurations, and traffic scenarios, models trained on one dataset may not perform well on another.
-For instance, CP models trained in simulation environments often struggle when applied to real-world scenarios.
-Future research should explore unsupervised domain adaptation and self-supervised learning methods to enhance cross-domain generalization.
-Additionally, using Generative Adversarial Networks (GANs) to synthesize cross-domain data and leveraging large-scale pre-trained models for domain transfer learning could provide effective solutions.
+Domain adaptation is a core challenge for improving model generalization across datasets.
+Differences in sensors, environments, and traffic patterns often degrade performance when models are transferred across domains.
+In particular, CP models trained on simulation data struggle in real-world deployment due to noise, occlusion, and latency mismatches.
+Future research should explore unsupervised and self-supervised methods to bridge this Sim2Real gap.
+Promising directions include domain-aligned feature learning, simulation-aware augmentation, and leveraging large-scale pre-trained models.
+Additionally, Generative Adversarial Networks can help synthesize realistic cross-domain samples to support robust transfer.
 VIII-E Privacy and Security
@@
 IX Conclusion
-This review presents a comprehensive review of CP datasets for AD, classifying them by collaboration paradigms, sensor modalities, data collection methods, environmental scenarios, and supported tasks.
-We analyze both real-world and simulation-based datasets across V2V, V2I, V2X, and I2I settings, evaluating their annotation quality, accessibility, scenario coverage, and benchmarking readiness.
-While existing datasets have significantly advanced CP research, limitations remain in data scale, coverage of rare and adverse conditions, agent diversity, and geographic generalization.
-To bridge these gaps, future datasets should incorporate long-tail scenarios such as traffic anomalies and extreme weather, include data from diverse regions, support heterogeneous agent types, and reflect realistic communication constraints.
-Establishing standardized annotation formats and unified evaluation protocols is essential for enabling fair comparisons and fostering cross-dataset generalization.
-Additionally, techniques such as automated labeling, high-fidelity simulation, and generative modeling—especially with the aid of large language models—offer promising directions for building scalable and semantically rich datasets.
-We hope this review provides valuable guidance for the design, construction, and benchmarking of future datasets, ultimately advancing CP systems in real-world AD applications.
+This review provides a comprehensive overview of CP datasets for AD, categorized by collaboration paradigms, sensor modalities, data sources, environments, and supported tasks.
+We examine real and simulated datasets across V2V, V2I, V2X, and I2I settings, assessing their annotation quality, accessibility, scenario diversity, and benchmarking readiness.
+Despite notable progress, challenges remain in data scale, rare condition coverage, agent diversity, and regional generalization.
+Future datasets should address long-tail scenarios, broaden geographic diversity, support heterogeneous agents, and capture realistic communication constraints.
+Standardized annotations and unified evaluation protocols are critical for fair comparison and cross-dataset generalization.
+Advances such as automated labeling, high-fidelity simulation, and LLM-assisted generative modeling offer promising pathways for scalable, semantically rich datasets.
+We hope this review informs future dataset design and accelerates CP research toward robust real-world AD applications.